### v1

# Towards Goal-oriented Large Language Model Prompting: A Survey

Abstract
Large Language Models (LLMs) have shown prominent performance in various downstream tasks in which prompt engineering plays a pivotal role in optimizing LLMs’ performance. This paper, not as an overview of current prompt engineering methods, aims to highlight the limitation of designing prompts while holding an anthropomorphic assumption that expects LLMs to think like humans. From our review of 35 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs. Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework by summarizing ten applicable tasks. With four future directions proposed, we hope to further emphasize and promote goal-oriented prompt engineering.
1 Introduction
Large Language Models (LLMs) like ChatGPT have garnered significant interest for their ability to assimilate vast amounts of information from large corpora of data. They have demonstrated proficiency in, for example, open-domain dialogue Deng et al. (2023); Yu et al. (2023), reasoning Wei et al. (2022); Yao et al. (2023); Besta et al. (2023), planning Song et al. (2023); Liu et al. (2023a), and so on. The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into their potential Liu et al. (2023b). However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al. (2021). Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
To obtain an optimal prompt for tasks, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information. Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996). In interacting with LLMs, which often display human-like conversational abilities, we might be tempted to attribute human-like thought processes to them, expecting them to know from the start that they should decompose complex problems into simpler tasks. This anthropomorphic assumption does not hold for LLMs and can lead to unsatisfactory outcomes Abercrombie et al. (2023). On the other hand, as shown in our literature review of 35 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly. In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers generated from LLMs that result in more effective and reliable human-computer interaction.
In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action implementation, sub-goal result evaluation, and sub-goal selection. We summarize the application of reviewed works on a wide variety of tasks to showcase the broad impact of goal-oriented philosophy in prompt engineering. Finally, we offer four main promising future directions, including the synergy of stages in the framework, applications to other tasks, efficiency problems, and hierarchical decomposition, for further improvement.
2 Taxonomy of Methods
Figure 1: An overview of the goal-oriented framework for prompting LLMs taking solving a math word problem as an example. (1) Decomposing goal into sub-goal sequences. (2) Action selection for attaining sub-goals. (3) Implementing actions to get sub-goal results. (4) Evaluation of sub-goal achievement based on sub-goal results. (5) Selecting valuable sub-goals for effective goal achievement. Note that stages (2)(3)(4) are taken for all the decomposed sub-goals.
In this section, we categorize existing methods into a goal-oriented framework, as shown in Fig. 1. We further refine the works in each part of the framework according to their distinctive features.
2.1 Decomposing Goal into Sub-goal Sequences
Decomposing a high-level goal into sub-goals is particularly useful for complex problems where a straightforward answer isn’t sufficient. In this section, we introduce the methods to decompose goals and auxiliary techniques to help improve the performance of decomposition.
Iterative decomposition.
Iterative decomposition generates a sub-goal and executes actions to get subgoal results, then repeat this process with the knowledge of the previous sub-goal and sub-goal results. Chain-of-thought prompting (CoT) Wei et al. (2022) can be considered the first work to try to decompose a goal into sub-goal sequences. By showing LLMs a series of intermediate reasoning steps that lead to the final output in the prompt, LLMs can naturally imitate a human-like problem-solving process. Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to the final solution of the problem. Kojima et al. (2022) even found that simply adding “Let’s think step by step.” to the prompt can guide LLMs to perform CoT decomposition as well. The above two works implicitly follow iterative decomposition since LLMs generate tokens in an autoregressive way which means that they can decide the next sub-goal based on previous content.
There are also works that explicitly ask LLMs to follow the iterative decomposition. DecomP Khot et al. (2022) and Successive Prompting Dua et al. (2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks. The background knowledge attained by answering each sub-question are sub-goals to be achieved. Different from CoT which may generate sub-questions sequentially in one output, these two works explicitly repeat prompting LLMs to ask follow-up questions. Self-ask Press et al. (2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions. Thus, LLMs can generate all essential questions and their answers by prompting once. Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions. This finding indicates the significance of decomposing complex goals into simple sub-goals for LLMs.
Plan-then-execute decomposition.
In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence at once, which means that the latter sub-goal will not be affected by the former ones. For example, Least-to-most prompting Zhou et al. (2022) only prompts LLMs two times, one for generating a plan to decompose the goal into a sub-goal sequence, and the other one for executing the plan. Plan-and-solve prompting Wang et al. (2023a) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output. DEPS Wang et al. (2023c) and GITM Zhu et al. (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft many types of items and tools. In this game, obtaining base materials that are needed to craft a given item can be viewed as sub-goals. While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
External decomposition.
The above two categories rely on LLM’s knowledge to decompose the goal into sub-goals. However, due to their hallucination problem Ji et al. (2023), they sometimes generate seemingly plausible sub-goals but not grounded in reality. To ensure the accuracy of decomposition, LLM+P Liu et al. (2023a) and SayPlan Rana et al. (2023) take advantage of classical planners. They use LLMs to translate goals written in natural language to planning domain definition language so that classical planners can deal with them. The output of planners will then be translated back into natural language by LLMs for execution.
Reduce sub-goal space.
Choosing possible sub-goals from a limited space is beneficial for getting a more precise and efficient sub-goal sequence since it avoids LLM getting distracted by irrelevant or even wrong sub-goals. PEARL Sun et al. (2023) is specially designed for question answering over long documents. Several sub-goals like “Finding the definition of A”, “Compare A and B”, and “Summarize A” are pre-set and LLMs only need to choose and arrange valuable sub-goals into a plan. Similar to PEARL, ProCoT Deng et al. (2023) has pre-defined sub-goal sets covering query clarification, topic transition, and negotiation strategy for dialogue systems. DecomP Khot et al. (2022) selects sub-goals from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task. SayPlan Rana et al. (2023) is designed for robot planning. Given a task instruction, semantic search is conducted to identify a relevant subgraph from the whole 3D scene graph that serves as the environment. LLMs will plan based only on this subgraph.
2.2 Action Selection for Attaining Sub-goals
Action selection is vital for achieving sub-goals as it involves choosing the most effective and valid actions to reach desired outcomes. CoT is a naive way for action selection where it is completely decided by LLMs themselves. However, due to the hallucination problem which is the same in decomposing sub-goals, actions generated by LLMs are often invalid. In this section, apart from naive selection, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
Constrain-then-select.
These kinds of methods first set constraints on the action space and then let LLMs decide the actual action. It shares similar ideas with the methods aimed at reducing sub-goal space, which makes it easier for LLMs to select valid actions. MWP Zhang et al. (2023) is proposed for solving math word problems. To ensure the correctness and coherence of the intermediate solution steps, the authors employ an operation prediction module that predicts the needed calculation operation (e.g. multiplication, summation) in this step. LLMs are used to fill the values in this template. In dialogue systems, RLP Hao et al. (2023a) additionally sets a mental state to reflect its personality traits to guide action selection. The authors found that this strategy facilitates contextually rich, coherent, and engaging interactions for LLM-based conversational systems.
Select-then-mapping.
Different from constrain-then-select, select-then-mapping first uses LLMs to generate actions solely based on their knowledge and then maps the generated action to the most similar one in the valid action space. Zero-shot planners Huang et al. (2022a), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al. (2022) aim to solve agent planning problems in an interactive environment. In such virtual environments, there are only a few admissible actions that can be applied. However, the actions produced naively by LLMs often cannot be mapped exactly to those executable actions. To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity. When users have access to all the admissible actions, it is suitable to define a set for action mapping.
Reference for selection.
It is empirically evaluated by some works that reference actions are helpful for LLMs to decide on actual actions. Reference action means the successful actions LLMs took when experiencing similar situations in the past. Intuitively, taking reference is similar to human skill improvement through practice. GITM Zhu et al. (2023), created for Minecraft, designs a text-based memory mechanism for LLM to store and retrieve gained useful knowledge during action selection. SALP Gramopadhye and Szafir (2023) collects references by choosing examples that have similar tasks to the given query. The examples are added to the input to autoregressively prompt LLMs. Here we want to discriminate between the reference examples and the in-context examples. In-context examples, which are used in CoT, are used to help LLMs understand and perform tasks Brown et al. (2020). For example, if a prompt includes intermediate steps for a reasoning problem, models will understand that they should also provide intermediate steps. The selection of reference examples in the prompt is stricter than the in-context examples. They should not only be identical in format, which is the criteria for selecting in-context examples, but also be similar in content.
2.3 Implementing Actions to Get Sub-goal Results
In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result. Due to the hallucination problem, the results are sometimes wrong. For example, even if LLMs correctly choose multiplication to solve a math problem, the calculation result is just not correct. Besides, as Khot et al. (2022) claimed, some tasks may not be feasible to solve using only an LLM.
Some works are proposed to leverage external tools to guarantee a precise sub-goal result. When deployed to answer open-domain questions, DecomP Khot et al. (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia. Program-of-Thoughts Chen et al. (2022) first translates reasoning processes into Python codes and then runs them in an interpreter. Toolformer Schick et al. (2023) specifically studied which APIs to call and when to call them when using external tools. In their paper, several tools are adopted, including a calculator for mathematical calculation, a calendar for the awareness of time, a BM25-based Wikipedia search engine for getting comprehensive subject information, a fine-tuned language model (LM) for question answering, and a fine-tuned LM for machine translation. Note that it adopts smaller LMs for specific tasks because authors found that fine-tuned LMs outperform LLMs on some basic functionalities. Following the idea of Toolformer, HuggingGPT Shen et al. (2023) further extends the available LMs to the scale of all LMs on open-sourced machine learning communities like HuggingFace. Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks. Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
2.4 Evaluation of Sub-goal Achievement based on Sub-goal Results
Feedback Source Method
LLM Self-refine Madaan et al. (2023)
SelfCheck Miao et al. (2023)
Reflexion Shinn et al. (2023)
Environment SayPlan Rana et al. (2023)
Re-prompt Raman et al. (2022)
Inner Monologue Huang et al. (2022b)
GITM  Zhu et al. (2023)
DEPS  Wang et al. (2023c)
VLM DEPS  Wang et al. (2023c)
LLM-Planner Song et al. (2023)
Inner Monologue Huang et al. (2022b)
Program Executor Self-debug Chen et al. (2023)
Heuristic Rule Reflexion Shinn et al. (2023)
Human Inner Monologue Huang et al. (2022b)
DEPS  Wang et al. (2023c)
Table 1: Feedback source used by existing works.
In CoT, there is no evaluation process for the sub-goal results. The results generated by LLMs are assumed to be correct by default. Similar to action implementation and action selection, sub-goal results suffer from the hallucination problem as well. In addition, mistakes in achieving one sub-goal may accumulate, which can lead to huge deviations from the true answer. Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time. Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well. The types of feedback source and their related works are summarized in Table 1.
Evaluated by LLMs.
Self-refine Madaan et al. (2023) is the first work addressing the evaluation of sub-goal results. After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output. For example, the prompt for providing feedback on code optimization problems may address efficiency and readability. Then, the feedback and initial output are fed together to LLMs for output refinement. Getting feedback and refinement are implemented as an iterative process. During iterations, all of the past feedback and refinement are appended to the prompt. For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified. SelfCheck Miao et al. (2023) decomposes the Self-refine evaluation into finer-grained processes. First, actions and sub-goal results are put to LLMs to summarize their intention. Then, the goal combined with all previous sub-goals, selected actions, and sub-goal results serves as another input to let LLMs extract useful information from them. Finally, only being provided with useful information and summarized intention, LLMs generate the sub-goal result again. This result will be compared with the original result. If the regeneration result supports or contradicts the original sub-goal result, it is likely that the original sub-goal result is correct or incorrect, respectively. In Reflexion Shinn et al. (2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result. The score is considered as feedback.
Figure 2: Illustration of approaches for valuable sub-goal selection. (a) CoT selects all sub-goals in one sub-goal sequence. (b) Self-consistency selects sub-goals based on majority votes. (c) MCR selects sub-goals from multiple sub-goal sequences. (d) Recmind and Selection-Inference select one sub-goal from candidates at each step. (e) ToT explores sub-goals from a tree structure space. (f) GDP-Zero and RAP introduce backpropagation to ToT to balance exploration and exploitation. (g) GoT models all sub-goals in a graph structure space.
Evaluated by external evaluators.
Self-refine and SelfCheck aim to prompt LLMs to stimulate the ability of self-correction. Alternatively, we could seek external evaluators for help to provide feedback, which is similar to the idea of applying external tools for action implementation.
Self-debug Chen et al. (2023) leverages unit tests and executors to gain feedback on generated code snippets. The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection. This finding indicates that external evaluators may be more beneficial than self-reflection thanks to more precise feedback. SayPlan Rana et al. (2023) and Re-prompting Raman et al. (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions. These two methods rely on the error messages that are pre-built into the program of virtual environments. The quality of error messages (e.g. detail) has a significant influence and sometimes we can hardly define error messages in advance. LLM-Planner Song et al. (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed. Inner Monologue Huang et al. (2022b) is a combination of SayPlan and LLM-Planner. It not only gets error messages from environments but also leverages VLMs to attain context information. What’s more, the VLM can also perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information. Similar to Inner Monologue, DEPS Wang et al. (2023c), a method proposed for Minecraft, also gets feedback from both VLMs and environments. DEPS further introduces a step to ask LLMs to explain the reason according to the feedback. GITM Zhu et al. (2023), also used for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation. All of the information can be obtained by calling Minecraft APIs. Reflexion Shinn et al. (2023) proposes various evaluators to handle different tasks. It has reward functions for reasoning tasks based on exact match grading and heuristic functions tailored to specific criteria for decision-making tasks. We refer to them as heuristic rules. Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine. Instead of implementing the action to get the sub-goal result again, these methods restart from action selection.
Reference for evaluation.
Recall that LLM-Planner does not have feedback from the environment. And in Reflexion it can only get a score value indicating the quality of the sub-goal result. For such methods, reference examples play a crucial role in helping LLMs evaluate the sub-goal result with more detail Shinn et al. (2023). The reference selection criteria are the same as in action selection, and examples with similar actions are more beneficial Song et al. (2023). For example, “cook a potato” is likely more informative than “clean a plate” for a sub-goal result from “cooking an egg”. While Reflexion builds a small memory buffer to store recent trials, LLM-Planner selects reference examples from the whole training sets based on the Euclidean distance between the BERT embeddings of references and the running trial.
2.5 Selecting Valuable Sub-goals for Effective Goal Achievement
In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process. Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant. One of the solutions, described in Section 2.1, is to reduce sub-goal space before asking LLMs to decide sub-goals. We denote such methods as sub-goal preprocess. Alternatively, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions. Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement. We denote this selection process as sub-goal postprocess. In this section, we introduce methods for sub-goal postprocess, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al. (2023) and its variants, and finally Graph of thoughts (GoT) Besta et al. (2023). An illustration of these approaches is shown in Fig. 2.
Chain-of-thought and its variants.
As a naive way of prompting LLMs, CoT asks LLMs to decompose the goal by themselves. The sub-goal selection is implicitly incorporated into the autoregressive decoding process where tokens with higher probabilities are more likely to be generated. However, the selected sub-goals may be irrelevant or even wrong due to the hallucination problem, which leads to the failure of goal achievement.
Four works are proposed to improve the robustness of CoT through sub-goal postprocess. Self-consistency Wang et al. (2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of the final output to decide valuable sub-goal sequences. The result from the chosen sequence is considered the final result. MCR Yoran et al. (2023) improves Self-consistency by considering all of the sub-goals in various sequences instead of the sub-goal sequence result. After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and take the result from the meta path. Recmind Wang et al. (2023b) and Selection-Inference Creswell et al. (2022) expand the sub-goal to multiple candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
Tree of thoughts and its variants.
ToT Yao et al. (2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure. To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search or depth-first search. For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs. CoT and its variants can be considered special cases of ToT.
RAP Hao et al. (2023b) and GDP-Zero Yu et al. (2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS). Compared with ToT, MCTS has an additional step for backpropagation, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence. This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
Graph of thoughts.
GoT Besta et al. (2023) advances ToT by modeling all of the LLM-generated sub-goals as a graph. While in ToT, only sub-goals in a sequence or at the same step can interact with each other through comparison, lookahead, or backtracking, GoT allows the synergy of arbitrary sub-goals. Specifically, GoT supports sub-goal aggregation to generate a new sub-goal and sub-goal refinement based on other dependent sub-goals. Finally, valuable sub-goals are selected via LLM self-evaluation on the constructed graph.
3 Application
Application Sub-category Stage
Method
Reasoning Arithmetic Reasoning ①②③④⑤
CoT Wei et al. (2022), Self-consistency Wang et al. (2022), Least-to-most Zhou et al. (2022), PoT Chen et al. (2022), Self-refine Madaan et al. (2023), Plan-and-solve Wang et al. (2023a), MWP Zhang et al. (2023), RAP Hao et al. (2023b), SelfCheck Miao et al. (2023), Toolformer Schick et al. (2023), Successive Prompt Dua et al. (2022)
Commonsense Reasoning
CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023a), Toolformer Schick et al. (2023)
Symbolic Reasoning
Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023a)
Logical Reasoning
Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023b)
Planning Virtual Environment ①②④
Zero-shot Planner Huang et al. (2022a), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al. (2023b), Reflexion Shinn et al. (2023), LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023c), GITM Zhu et al. (2023), Inner Monologue Huang et al. (2022b)
Real Environment
Inner Monologue Huang et al. (2022b)
Question Answering Multihop QA ①③④⑤
DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
Open domain QA
Toolformer Schick et al. (2023)
Code Generation
Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023)
Dialogue ①④⑤
ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
Recommendation
Recmind Wang et al. (2023b)
Table 2: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1. We maintain a list of related datasets and miscellaneous tasks at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering.
In this section, we provide an overview of the application of goal-oriented prompt engineering on LLMs. Specifically, we first introduce the definition of the task and then describe how current methods, following into different stages within our framework, leverage goal-oriented prompt engineering on LLMs to achieve task objectives. A summary of the applicable tasks, stages involved, and respective research works are shown in Table 2. Note that here we only introduce representative tasks. For miscellaneous tasks and related datasets, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering to give more comprehensive information. We can see that our framework can be applied to a wide range of tasks.
Reasoning.
There are specifically four types of reasoning tasks: arithmetic reasoning, commonsense reasoning, logical reasoning, and symbolic reasoning. These reasoning tasks focus on math word problems, world knowledge, logical principles and rules, and manipulation of symbols such as letters, respectively. Reasoning involves multiple intermediate steps to obtain a final answer. Additionally, to obtain the correct answer the LLM must interpret the question correctly, extract necessary information from the given context, and maintain consistency within its intermediate steps.
The goal of these tasks can be considered as producing an answer to a given question. The intermediate steps can be viewed as sub-goals that are sequentially connected to reach the final goal Wei et al. (2022); Wang et al. (2023a); Khot et al. (2022). The sub-goal results from action implementation, which in this case are intermediate conclusions, are often evaluated and revised to avoid error accumulation which may negatively influence the following derivations Madaan et al. (2023); Miao et al. (2023). Alternatively, for certain actions like calculations, we could leverage an external calculator to ensure a precise sub-goal result Schick et al. (2023) and employ sketches to help equation formulation (action selection) Zhang et al. (2023). We can also ask LLMs to explore parallel possible intermediate conclusions to form an overall better solution.
Planning.
Given an objective like “Fetch a bottle from the kitchen”, planning involves developing a series of sub-objectives (e.g. “Go to the kitchen”, “Find the cupboard”) Liu et al. (2023a); Song et al. (2023) and reaching sub-objectives by selecting appropriate actions (e.g. “Move”, “Search”) Hao et al. (2023a); Raman et al. (2022) to achieve the overall goal in a virtual or real environment.
We can see that works focusing on this task do not propose methods for action implementation and valuable sub-goal selection. This is because the evaluated environments are relatively straightforward and existing works mainly focus on the accuracy of planning without considering efficiency. LLMs often interact with virtual environments through APIs Wang et al. (2023c); Zhu et al. (2023) or real environments through mechanical devices Huang et al. (2022b) to ensure the actions are implemented precisely.
Question Answering.
Open domain question answering is a task aimed at answering factual questions without any explicit evidence. Multihop QA is an advanced task over open-domain QA. The former relies more on world knowledge while the latter requires multi-step reasoning based on world knowledge to answer more complex questions.
Considering answering complex questions as the goal, LLMs are asked to first decompose them into basic sub-questions Khot et al. (2022); Sun et al. (2023), gain knowledge for sub-questions Schick et al. (2023), and reflect on the correctness of the knowledge Press et al. (2022); Shinn et al. (2023). Sometimes LLMs may ask irrelevant sub-questions or give wrong answers to those sub-questions. To avoid their negative influence, there are methods proposed to lead LLMs to select valuable sub-questions from all candidates Yoran et al. (2023).
Code Generation.
Code generation aims to write a code snippet according to the given description. In this task, our goal is to generate code that is both functionally correct and aligns with the description. All of the works applied to this task fall into the evaluation of results. They leverage the error messages from unit tests Shinn et al. (2023); Chen et al. (2023) or rely on LLMs themselves Chen et al. (2023); Madaan et al. (2023) to reflect the validity of generated code snippets and revise them if needed.
Dialogue.
Dialogue systems are designed to provide functional service or social support to users via interactions. LLMs often play the role of a conversational agent to achieve certain goals (e.g. query clarification, offer emotional support). To give users positive interaction, LLMs have to correctly understand the context and generate clear and brief responses. Thus, some sub-goals are pre-defined according to the goal to ensure effective decomposition Deng et al. (2023). And LLMs are asked to reflect on their response to generate better content Madaan et al. (2023).
Recommendation.
The recommender system, which targets providing personalized recommendations to users based on their preferences, is crucial in many application scenarios including social media, e-commerce websites, etc. Considering giving ratings for items as the final goal, collecting rating information from various sources are then sub-goals. To rely on more valuable sources, LLMs can be asked to compare their credibility and authenticity, and finally take the most valuable source as the overall ratings Wang et al. (2023b).
4 Challenges & Opportunities
In this section, we discuss the challenges and opportunities of goal-oriented prompt engineering. Even if there have been numerous works proposed to facilitate processing at each stage, there remain some potential directions.
Synergy of stages.
The synergy of each stage in our framework will further improve the performance of goal-oriented prompt engineering. We find that though numerous works are proposed to address different stages in the framework, no one synergizes them all.
For reasoning tasks, each stage has specially proposed methods, and putting them together is straightforward. However, for some tasks like code generation, the vacancy of works for certain stages hinders the integration. According to Table 2, there are only works focusing on sub-goal evaluation in code generation. To bridge the gap, we have to understand how the code generation process can fit into our framework. In fact, programming is a process of breaking down a problem into small, manageable parts and solving them in a logical sequence. For example, suppose we are tasked with calculating the average of a set of numbers. We can break this problem down into the following steps: (1) Input a set of numbers. (2) Calculate the sum of these numbers. (3) Determine the count of numbers. (4) Use the sum and count to calculate the average. Considering these intermediate steps as sub-goals, we decompose the goal of finishing a programming problem into a sequence of sub-goals. For other stages like action selection and sub-goal evaluation, we could apply methods that help LLMs efficiently select actions to calculate the summation or reflect whether their intermediate answer is correct.
Application to other tasks.
Though existing works have covered up to 11 tasks, goal-oriented prompt engineering has the potential for broader applications. For example, apart from implementing functions, code review also plays a crucial role in software development. Various review criteria can be taken as sub-goals to ensure the quality of the code, which is our overall goal. We may prompt LLMs to design approaches to evaluate different criteria (e.g. text-based similarity measurement for coding style evaluation, unit tests generation for functionality verification), which corresponds to the action selection to achieve sub-goals. Similar to Program-of-thoughts Chen et al. (2022), we can use external interpreters to test the code. For some criteria without precise success or failure signals, we can apply sub-goal self-evaluation methods to facilitate better feedback. This fitting-in is also suitable for requirement analysis where we also have to think through several perspectives. Taking one step further, we think this framework even has the potential to be applied to the whole life cycle of software development.
Efficiency.
The works we review in this paper mainly address the accuracy of goal-oriented prompt engineering. In other words, they care whether the goal is achieved finally. However, efficiency also plays a crucial role in application. As we are optimistic about the accuracy when employing methods in each stage of our framework to prompt design, we think efficiency is the next key factor that affects performance. For example, in planning tasks, LLMs are expected to form a plan that costs the least energy of human beings or robots. And in dialogue systems, users prefer to achieve their goals with as few interaction rounds as possible. To evaluate efficiency, we need to either design heuristic rules for automatic measurement or annotate the resource-consuming information by human beings.
Hierarchical decomposition.
Existing works decompose the goal into a sub-goal sequence, which can be limiting. The method might fall short in adequately addressing the intricacies of each sub-goal, especially in situations where each sub-goal itself is complex and multifaceted. Thus, LLMs may not achieve certain sub-goals perfectly, then the error will be accumulated, leading to the failure of overall goal achievement. To make sure each sub-goal is simple enough to be handled by LLMs, we introduce the idea of hierarchical decomposition where sub-goals can be decomposed into even more simple sub-goals. GITM Zhu et al. (2023) could be viewed as employing the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones. Though a significant success rate increase by employing hierarchical decomposition is reported, relying on solely a tree may not be able to handle more complex problems, which leaves room for further improvement.

### v2

@@
-Towards Goal-oriented Large Language Model Prompting: A Survey
+Towards Goal-oriented Prompt Engineering for Large Language Models: A Survey
 Abstract
-Large Language Models (LLMs) have shown prominent performance in various downstream tasks in which prompt engineering plays a pivotal role in optimizing LLMs’ performance.
-This paper, not as an overview of current prompt engineering methods, aims to highlight the limitation of designing prompts while holding an anthropomorphic assumption that expects LLMs to think like humans.
-From our review of 35 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs.
-Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework by summarizing ten applicable tasks.
-With four future directions proposed, we hope to further emphasize and promote goal-oriented prompt engineering.
+Large Language Models (LLMs) have shown prominent performance in various downstream tasks and prompt engineering plays a pivotal role in optimizing LLMs’ performance.
+This paper, not only as an overview of current prompt engineering methods, but also aims to highlight the limitation of designing prompts based on an anthropomorphic assumption that expects LLMs to think like humans.
+From our review of 36 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs.
+Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework.
+With four future directions proposed, we hope to further emphasize the power and potential of goal-oriented prompt engineering in all fields.
+Towards Goal-oriented Prompt Engineering
+for Large Language Models: A Survey
+Haochen Li    Jonathan Leung    Zhiqi Shen Nanyang Technological University, Singapore {haochen003, jonathan008, zqshen}@ntu.edu.sg
 1 Introduction
-Large Language Models (LLMs) like ChatGPT have garnered significant interest for their ability to assimilate vast amounts of information from large corpora of data.
-They have demonstrated proficiency in, for example, open-domain dialogue Deng et al.
+Large Language Models (LLMs) have garnered significant interest with their ability to assimilate vast amounts of information from large corpora of data.
+They have demonstrated proficiency in open-domain dialogue Deng et al.
 (2023); Yu et al.
@@
 (2023); Liu et al.
-(2023a), and so on.
-The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into their potential Liu et al.
-(2023b).
+(2023a), and many more tasks.
+The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into LLMs’ potential Liu et al.
+(2023c).
 However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al.
@@
 Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
-To obtain an optimal prompt for tasks, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information.
+To prompt LLMs appropriately, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information.
 Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996).
@@
 (2023).
-On the other hand, as shown in our literature review of 35 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly.
-In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers generated from LLMs that result in more effective and reliable human-computer interaction.
-In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action implementation, sub-goal result evaluation, and sub-goal selection.
-We summarize the application of reviewed works on a wide variety of tasks to showcase the broad impact of goal-oriented philosophy in prompt engineering.
-Finally, we offer four main promising future directions, including the synergy of stages in the framework, applications to other tasks, efficiency problems, and hierarchical decomposition, for further improvement.
-2 Taxonomy of Methods
+On the other hand, as shown in our literature review of 36 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly.
+In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers from LLMs and results in more effective and reliable human-computer interaction.
 Figure 1: An overview of the goal-oriented framework for prompting LLMs taking solving a math word problem as an example.
@@
 (2) Action selection for attaining sub-goals.
-(3) Implementing actions to get sub-goal results.
-(4) Evaluation of sub-goal achievement based on sub-goal results.
-(5) Selecting valuable sub-goals for effective goal achievement.
+(3) Executing actions to get sub-goal results.
+(4) Evaluating sub-goal results.
+(5) Further selection of valuable sub-goals.
 Note that stages (2)(3)(4) are taken for all the decomposed sub-goals.
-In this section, we categorize existing methods into a goal-oriented framework, as shown in Fig. 1. We further refine the works in each part of the framework according to their distinctive features.
+While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures.
+Some surveys summarize studies that focus on specific tasks or problems.
+Qiao et al.
+(2022); Huang and Chang (2022) focus solely on reasoning tasks and Wang et al.
+(2024) on planning tasks.
+Huang et al.
+(2023) deals with the hallucination problem of LLMs.
+Other studies focus on generally applicable prompting but different sub-fields.
+Amatriain (2024); Sahoo et al.
+(2024) simply listed out recent techniques without systematic categorization.
+Liu et al.
+(2023c) focus on the prompting framework for facilitating the interaction between LLMs with external worlds.
+Liu et al.
+(2023b) focuses on methods for learnable continuous prompt, while our work studies discrete prompt.
+As many of the will-functioned LLMs are not open-sourced (e.g.
+GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
+In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action execution, sub-goal result evaluation, and sub-goal selection, as shown in Fig.
+1.
+We further classify the surveyed methods by their targeted task and analyze their respective performance to provide readers with a better estimate of the current progress of goal-oriented prompt engineering.
+Finally, we offer four promising future directions, including hierarchical decomposition, efficiency problems, the synergy of stages, and broader applications.
+2 Taxonomy of Methods
+In this section, we go into detail about the five steps described in our goal-oriented framework in Fig. 1.
 2.1 Decomposing Goal into Sub-goal Sequences
 Decomposing a high-level goal into sub-goals is particularly useful for complex problems where a straightforward answer isn’t sufficient.
-In this section, we introduce the methods to decompose goals and auxiliary techniques to help improve the performance of decomposition.
+In this section, we introduce the methods to decompose goals and improve the performance of decomposition.
 Iterative decomposition.
-Iterative decomposition generates a sub-goal and executes actions to get subgoal results, then repeat this process with the knowledge of the previous sub-goal and sub-goal results.
+Iterative decomposition generates a sub-goal, gets sub-goal results, and repeats this process with the knowledge of the previous sub-goal and sub-goal results.
 Chain-of-thought prompting (CoT) Wei et al.
-(2022) can be considered the first work to try to decompose a goal into sub-goal sequences.
-By showing LLMs a series of intermediate reasoning steps that lead to the final output in the prompt, LLMs can naturally imitate a human-like problem-solving process.
-Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to the final solution of the problem.
-Kojima et al.
-(2022) even found that simply adding “Let’s think step by step.” to the prompt can guide LLMs to perform CoT decomposition as well.
-The above two works implicitly follow iterative decomposition since LLMs generate tokens in an autoregressive way which means that they can decide the next sub-goal based on previous content.
+(2022) can be considered the first work that follows iterative decomposition.
+With either “Let’s think step by step.” Kojima et al.
+(2022), or intermediate reasoning steps in in-context examples in the prompt, LLMs can imitate the decomposition process and solve the problem step by step.
+Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to a correct output.
+CoT implicitly follows iterative decomposition as LLMs generate tokens in an autoregressive way, with the next sub-goal decided based on previous content.
 There are also works that explicitly ask LLMs to follow the iterative decomposition.
@@
 (2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks.
-The background knowledge attained by answering each sub-question are sub-goals to be achieved.
-Different from CoT which may generate sub-questions sequentially in one output, these two works explicitly repeat prompting LLMs to ask follow-up questions.
+The background knowledge required to answer the question is sub-goal to be achieved.
+Different from CoT which may generate sub-questions sequentially in one output, these two works repeatedly prompt LLMs to ask follow-up questions.
 Self-ask Press et al.
 (2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions.
-Thus, LLMs can generate all essential questions and their answers by prompting once.
+With that template, LLMs can generate all essential questions and their answers by prompting once.
 Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions.
@@
 Plan-then-execute decomposition.
-In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence at once, which means that the latter sub-goal will not be affected by the former ones.
+In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence all at once, which means that the latter sub-goal will not be affected by the result of the former ones.
 For example, Least-to-most prompting Zhou et al.
@@
 Plan-and-solve prompting Wang et al.
-(2023a) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output.
+(2023b) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output.
 DEPS Wang et al.
-(2023c) and GITM Zhu et al.
-(2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft many types of items and tools.
-In this game, obtaining base materials that are needed to craft a given item can be viewed as sub-goals.
+(2023d) and GITM Zhu et al.
+(2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft different items and tools.
+In this game, obtaining base materials that are needed to craft a target item can be viewed as sub-goals.
 While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
@@
 The above two categories rely on LLM’s knowledge to decompose the goal into sub-goals.
-However, due to their hallucination problem Ji et al.
-(2023), they sometimes generate seemingly plausible sub-goals but not grounded in reality.
+However, due to the hallucination problem Ji et al.
+(2023), LLMs sometimes generate seemingly plausible sub-goals but not grounded in reality.
 To ensure the accuracy of decomposition, LLM+P Liu et al.
@@
 The output of planners will then be translated back into natural language by LLMs for execution.
-Reduce sub-goal space.
-Choosing possible sub-goals from a limited space is beneficial for getting a more precise and efficient sub-goal sequence since it avoids LLM getting distracted by irrelevant or even wrong sub-goals.
+2.2 Action Selection for Attaining Sub-goals
+With sub-goals defined, another important step is action selection, choosing the most effective and valid actions to reach desired outcomes.
+CoT is a naive way for action selection where it is completely decided by LLMs themselves.
+However, due to the hallucination problem, actions generated by LLMs are often invalid.
+In this section, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
+Constrain-then-select.
+Constrain-then-select method predefines an action space and then has LLMs select the actual action among the space.
+As in MWP Zhang et al.
+(2023), to deal with math word problems, the authors first employ an operation prediction module to predict the needed calculation operation (e.g.
+multiplication, summation).
+LLMs are then used to select the appropriate operands and complete the calculation.
 PEARL Sun et al.
-(2023) is specially designed for question answering over long documents.
-Several sub-goals like “Finding the definition of A”, “Compare A and B”, and “Summarize A” are pre-set and LLMs only need to choose and arrange valuable sub-goals into a plan.
+(2023) utilizes constrained action space for question answering over long documents.
+The authors predefine actions like “Finding the definition of A”, “Compare A and B”, and “Summarize A”.
+DecomP Khot et al.
+(2022) selects actions from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task.
+SayPlan Rana et al.
+(2023) is for robot planning.
+Given a task instruction, the authors apply semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, then have LLMs plan based on the subgraph.
+In dialogue systems, RLP Fischer (2023) allows the user to set a mental state of LLM, such as beliefs or desires, to instruct action selection.
+Such a setting improves the context richness, coherence, and interactivity of LLM-based conversational systems.
 Similar to PEARL, ProCoT Deng et al.
-(2023) has pre-defined sub-goal sets covering query clarification, topic transition, and negotiation strategy for dialogue systems.
-DecomP Khot et al.
-(2022) selects sub-goals from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task.
-SayPlan Rana et al.
-(2023) is designed for robot planning.
-Given a task instruction, semantic search is conducted to identify a relevant subgraph from the whole 3D scene graph that serves as the environment.
-LLMs will plan based only on this subgraph.
-2.2 Action Selection for Attaining Sub-goals
-Action selection is vital for achieving sub-goals as it involves choosing the most effective and valid actions to reach desired outcomes.
-CoT is a naive way for action selection where it is completely decided by LLMs themselves.
-However, due to the hallucination problem which is the same in decomposing sub-goals, actions generated by LLMs are often invalid.
-In this section, apart from naive selection, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
-Constrain-then-select.
-These kinds of methods first set constraints on the action space and then let LLMs decide the actual action.
-It shares similar ideas with the methods aimed at reducing sub-goal space, which makes it easier for LLMs to select valid actions.
-MWP Zhang et al.
-(2023) is proposed for solving math word problems.
-To ensure the correctness and coherence of the intermediate solution steps, the authors employ an operation prediction module that predicts the needed calculation operation (e.g.
-multiplication, summation) in this step.
-LLMs are used to fill the values in this template.
-In dialogue systems, RLP Hao et al.
-(2023a) additionally sets a mental state to reflect its personality traits to guide action selection.
-The authors found that this strategy facilitates contextually rich, coherent, and engaging interactions for LLM-based conversational systems.
+(2023) forms pre-defined action sets covering query clarification, topic transition, and negotiation strategy for dialogue systems.
 Select-then-mapping.
@@
 To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity.
-When users have access to all the admissible actions, it is suitable to define a set for action mapping.
-Reference for selection.
-It is empirically evaluated by some works that reference actions are helpful for LLMs to decide on actual actions.
-Reference action means the successful actions LLMs took when experiencing similar situations in the past.
-Intuitively, taking reference is similar to human skill improvement through practice.
-GITM Zhu et al.
-(2023), created for Minecraft, designs a text-based memory mechanism for LLM to store and retrieve gained useful knowledge during action selection.
-SALP Gramopadhye and Szafir (2023) collects references by choosing examples that have similar tasks to the given query.
-The examples are added to the input to autoregressively prompt LLMs.
-Here we want to discriminate between the reference examples and the in-context examples.
-In-context examples, which are used in CoT, are used to help LLMs understand and perform tasks Brown et al.
-(2020).
-For example, if a prompt includes intermediate steps for a reasoning problem, models will understand that they should also provide intermediate steps.
-The selection of reference examples in the prompt is stricter than the in-context examples.
-They should not only be identical in format, which is the criteria for selecting in-context examples, but also be similar in content.
-2.3 Implementing Actions to Get Sub-goal Results
+2.3 Executing Actions to Get Sub-goal Results
 In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result.
 Due to the hallucination problem, the results are sometimes wrong.
-For example, even if LLMs correctly choose multiplication to solve a math problem, the calculation result is just not correct.
+For example, even if LLMs correctly select multiplication as the action to solve a math problem, the calculation result can be wrong.
 Besides, as Khot et al.
-(2022) claimed, some tasks may not be feasible to solve using only an LLM.
-Some works are proposed to leverage external tools to guarantee a precise sub-goal result.
-When deployed to answer open-domain questions, DecomP Khot et al.
+(2022) claimed, some tasks may not be feasible to solve using only one LLM.
+With such consideration, some works leverage external tools to guarantee a precise sub-goal result.
+Some studies specifically study one task Khot et al.
+(2022); Chen et al.
+(2022); Wang et al.
+(2023c).
+To answer open-domain questions, DecomP Khot et al.
 (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia.
-Program-of-Thoughts Chen et al.
+Program-of-Thoughts (PoT) Chen et al.
 (2022) first translates reasoning processes into Python codes and then runs them in an interpreter.
+Recmind Wang et al.
+(2023c) uses SQL to search for the history of a user in the database in recommender systems.
+There are additional studies that aim for multi-tasking.
 Toolformer Schick et al.
 (2023) specifically studied which APIs to call and when to call them when using external tools.
-In their paper, several tools are adopted, including a calculator for mathematical calculation, a calendar for the awareness of time, a BM25-based Wikipedia search engine for getting comprehensive subject information, a fine-tuned language model (LM) for question answering, and a fine-tuned LM for machine translation.
-Note that it adopts smaller LMs for specific tasks because authors found that fine-tuned LMs outperform LLMs on some basic functionalities.
+In their paper, several tools are adopted, including a calculator, a calendar, a BM25-based Wikipedia search engine, and a fine-tuned language model (LM).
+The authors pointed out that small LMs finetuned for specific tasks can outperform some LLMs on certain task.
 Following the idea of Toolformer, HuggingGPT Shen et al.
@@
 Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
-2.4 Evaluation of Sub-goal Achievement based on Sub-goal Results
+2.4 Evaluating Sub-goal Results
 Feedback Source Method
@@
 GITM  Zhu et al. (2023)
-DEPS  Wang et al. (2023c)
-VLM DEPS  Wang et al. (2023c)
+DEPS  Wang et al. (2023d)
+VLM DEPS  Wang et al. (2023d)
 LLM-Planner Song et al. (2023)
@@
 Program Executor Self-debug Chen et al. (2023)
+INTERVENOR Wang et al. (2023a)
 Heuristic Rule Reflexion Shinn et al. (2023)
 Human Inner Monologue Huang et al. (2022b)
-DEPS  Wang et al. (2023c)
+DEPS  Wang et al. (2023d)
 Table 1: Feedback source used by existing works.
@@
 The results generated by LLMs are assumed to be correct by default.
-Similar to action implementation and action selection, sub-goal results suffer from the hallucination problem as well.
-In addition, mistakes in achieving one sub-goal may accumulate, which can lead to huge deviations from the true answer.
+Answering by LLMs, sub-goal results also suffer from the hallucination problem.
+In addition, errors may propagate along the pipeline and result in huge deviations from the correct answer.
 Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time.
@@
 The types of feedback source and their related works are summarized in Table 1.
-Evaluated by LLMs.
+Figure 2: Illustration of approaches for valuable sub-goal selection.
+(a) CoT selects all sub-goals in one sub-goal sequence.
+(b) Self-consistency, a variant of CoT, selects sub-goals based on majority votes.
+(c) MCR, a variant of CoT, selects sub-goals from multiple sub-goal sequences.
+(d) Recmind and Selection-Inference, variants of CoT, select one sub-goal from candidates at each step.
+(e) ToT explores sub-goals from a tree structure space.
+(f) GDP-Zero and RAP, variants of ToT, introduce backpropagation to ToT to balance exploration and exploitation.
+(g) GoT models all sub-goals in a graph structure space.
+Evaluate with LLMs.
 Self-refine Madaan et al.
@@
 After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output.
-For example, the prompt for providing feedback on code optimization problems may address efficiency and readability.
 Then, the feedback and initial output are fed together to LLMs for output refinement.
-Getting feedback and refinement are implemented as an iterative process.
-During iterations, all of the past feedback and refinement are appended to the prompt.
+The refinement process iterates with all of the past feedback and refinement appended to the prompt.
 For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified.
@@
 This result will be compared with the original result.
-If the regeneration result supports or contradicts the original sub-goal result, it is likely that the original sub-goal result is correct or incorrect, respectively.
+If the regeneration result supports the original sub-goal result, the original result is considered correct, and vice versa.
 In Reflexion Shinn et al.
 (2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result.
-The score is considered as feedback.
-Figure 2: Illustration of approaches for valuable sub-goal selection.
-(a) CoT selects all sub-goals in one sub-goal sequence.
-(b) Self-consistency selects sub-goals based on majority votes.
-(c) MCR selects sub-goals from multiple sub-goal sequences.
-(d) Recmind and Selection-Inference select one sub-goal from candidates at each step.
-(e) ToT explores sub-goals from a tree structure space.
-(f) GDP-Zero and RAP introduce backpropagation to ToT to balance exploration and exploitation.
-(g) GoT models all sub-goals in a graph structure space.
-Evaluated by external evaluators.
-Self-refine and SelfCheck aim to prompt LLMs to stimulate the ability of self-correction.
-Alternatively, we could seek external evaluators for help to provide feedback, which is similar to the idea of applying external tools for action implementation.
+Evaluate with external evaluators.
+In contrast to self-evaluation, we could seek the help of external evaluators for providing feedback, which is similar to but different from the idea of applying external tools for action execution.
+Some methods Chen et al.
+(2023); Wang et al.
+(2023a); Rana et al.
+(2023); Raman et al.
+(2022) leverage the error messages that are pre-built into either the programming languages or the program of virtual environments to evaluate sub-goal results.
 Self-debug Chen et al.
-(2023) leverages unit tests and executors to gain feedback on generated code snippets.
+(2023) and INTERVENOR Wang et al.
+(2023a) leverages unit tests and executors to gain feedback on generated code snippets.
 The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection.
-This finding indicates that external evaluators may be more beneficial than self-reflection thanks to more precise feedback.
+This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback.
 SayPlan Rana et al.
@@
 (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions.
-These two methods rely on the error messages that are pre-built into the program of virtual environments.
+GITM Zhu et al.
+(2023), a method proposed for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation.
+All of the information can be obtained from Minecraft API returns.
 The quality of error messages (e.g.
-detail) has a significant influence and sometimes we can hardly define error messages in advance.
-LLM-Planner Song et al.
+detail) is influential to the evaluating process, but sometimes it is hard to define error messages in advance.
+To further address it, LLM-Planner Song et al.
 (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed.
 Inner Monologue Huang et al.
-(2022b) is a combination of SayPlan and LLM-Planner.
-It not only gets error messages from environments but also leverages VLMs to attain context information.
-What’s more, the VLM can also perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information.
+(2022b) is a combination of SayPlan and LLM-Planner, where VLMs additionally perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information.
 Similar to Inner Monologue, DEPS Wang et al.
-(2023c), a method proposed for Minecraft, also gets feedback from both VLMs and environments.
+(2023d), designed for Minecraft, also gets feedback from both VLMs and environments.
 DEPS further introduces a step to ask LLMs to explain the reason according to the feedback.
-GITM Zhu et al.
-(2023), also used for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation.
-All of the information can be obtained by calling Minecraft APIs.
 Reflexion Shinn et al.
@@
 Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine.
-Instead of implementing the action to get the sub-goal result again, these methods restart from action selection.
-Reference for evaluation.
-Recall that LLM-Planner does not have feedback from the environment.
-And in Reflexion it can only get a score value indicating the quality of the sub-goal result.
-For such methods, reference examples play a crucial role in helping LLMs evaluate the sub-goal result with more detail Shinn et al.
-(2023).
-The reference selection criteria are the same as in action selection, and examples with similar actions are more beneficial Song et al.
-(2023).
-For example, “cook a potato” is likely more informative than “clean a plate” for a sub-goal result from “cooking an egg”.
-While Reflexion builds a small memory buffer to store recent trials, LLM-Planner selects reference examples from the whole training sets based on the Euclidean distance between the BERT embeddings of references and the running trial.
-2.5 Selecting Valuable Sub-goals for Effective Goal Achievement
+Instead of executing the action to get the sub-goal result again, these methods restart from action selection.
+2.5 Further Selection of Valuable Sub-goals
 In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process.
 Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant.
-One of the solutions, described in Section 2.1, is to reduce sub-goal space before asking LLMs to decide sub-goals.
-We denote such methods as sub-goal preprocess.
-Alternatively, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions.
+To address this problem, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions.
 Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement.
-We denote this selection process as sub-goal postprocess.
-In this section, we introduce methods for sub-goal postprocess, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al.
+In this section, we introduce methods for such sub-goal postprocessing, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al.
 (2023) and its variants, and finally Graph of thoughts (GoT) Besta et al.
@@
 An illustration of these approaches is shown in Fig.
-2.
+2 to better distinguish these methods.
 Chain-of-thought and its variants.
-As a naive way of prompting LLMs, CoT asks LLMs to decompose the goal by themselves.
-The sub-goal selection is implicitly incorporated into the autoregressive decoding process where tokens with higher probabilities are more likely to be generated.
-However, the selected sub-goals may be irrelevant or even wrong due to the hallucination problem, which leads to the failure of goal achievement.
+We consider CoT as simply selecting all sub-goals since LLMs have access to all of the sub-goals during the generation process, which makes it prone to irrelevant sub-goals or wrong sub-goal results.
 Four works are proposed to improve the robustness of CoT through sub-goal postprocess.
 Self-consistency Wang et al.
-(2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of the final output to decide valuable sub-goal sequences.
+(2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences.
 The result from the chosen sequence is considered the final result.
 MCR Yoran et al.
-(2023) improves Self-consistency by considering all of the sub-goals in various sequences instead of the sub-goal sequence result.
-After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and take the result from the meta path.
+(2023) improves Self-consistency by selecting and combining sub-goals from different sequences.
+After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and predict the result based on the meta path.
 Recmind Wang et al.
-(2023b) and Selection-Inference Creswell et al.
-(2022) expand the sub-goal to multiple candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
+(2023c) and Selection-Inference Creswell et al.
+(2022) allow multiple sub-goal candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
 Tree of thoughts and its variants.
@@
 (2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure.
-To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search or depth-first search.
+To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search.
 For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs.
-CoT and its variants can be considered special cases of ToT.
+CoT and its variants can be considered special cases of ToT, where each node only has one child node.
 RAP Hao et al.
-(2023b) and GDP-Zero Yu et al.
+(2023) and GDP-Zero Yu et al.
 (2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS).
-Compared with ToT, MCTS has an additional step for backpropagation, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence.
+Compared with ToT, MCTS has an additional back-propagation step, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence.
 This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
@@
 While in ToT, only sub-goals in a sequence or at the same step can interact with each other through comparison, lookahead, or backtracking, GoT allows the synergy of arbitrary sub-goals.
-Specifically, GoT supports sub-goal aggregation to generate a new sub-goal and sub-goal refinement based on other dependent sub-goals.
+What’s more, GoT supports sub-goal aggregation to generate a new sub-goal and sub-goal refinement based on other dependent sub-goals.
 Finally, valuable sub-goals are selected via LLM self-evaluation on the constructed graph.
-3 Application
-Application Sub-category Stage
-Method
-Reasoning Arithmetic Reasoning ①②③④⑤
-CoT Wei et al.
+3 Discussion
+Application.
+After understanding how each method leverages goal-oriented prompt engineering to achieve goals with LLMs, we find it insightful to provide a bird-eye view of the surveyed paper.
+In Table 9, we classify these papers by their targeted tasks and the stages involved in our goal-oriented framework, to help readers better understand the current progress [^1] Note that here we only introduce representative tasks.
+For more comprehensive information, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering..
+The 36 representative works in our survey cover 11 tasks, including arithmetic reasoning (11 works), commonsense reasoning (5 works), symbolic reasoning (2 works), logical reasoning (3 works), planning in virtual environment (10 works), planning in real environment (1 work), multihop question answering (5 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (3 works), and recommender systems (1 work).
+We could see that 60% of the works focus on reasoning, 31% of the works are for planning and the remaining 9% of the works are for other tasks.
+For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 23% equally.
+There is potential for exploring the synergy of stages, which we will discuss in Section 4.
+For planning tasks, goal decomposition shares 31%, action selection shares 23%, and sub-goal result evaluation shares the rest 46%.
+No surveyed work for planning tasks involves the action execution stage and valuable sub-goal selection stage.
+This is because the currently available evaluation benchmarks are based on either simple real environments or carefully designed virtual environments.
+Actions are either easy or can be assured to be perfectly executed.
+Nonetheless, as the field develops, we can expect more complicated environments when actions cost differently or cannot be perfectly executed.
+Under such scenarios, there should be attention paid to action execution and valuable sub-goal selection stage.
+For those minority tasks like code generation and recommendation, most of the stages remain unexplored.
+With the goal-oriented framework proven effective in reasoning and planning tasks, it should be valuable to explore the potential of the framework on other tasks and we will discuss it in Sec 4.
+Performance.
+Furthermore, We compare the performance of the surveyed methods in Appendix A (Table 2, 3, 4, 5, 6, 7, 8).
+Compared with vanilla CoT, on average goal decomposition (Plan-and-solve) improves by 6.1%, action selection (PoT) improves by 2.9%, and valuable sub-goal selection (Self-consistency) improves by 22.7% in arithmetic reasoning, as shown in Table 6.
+Besides, we observe a gradual increase in performance when integrating additional stages into the method.
+When combined with Self-consistency, Plan-and-solve can be improved by 11.3% and PoT can be improved by 5.2% on average across several arithmetic reasoning datasets, which indicates that stage synergy may bring further improvement.
+During data collection, we find it hard to compare between methods due to the variety of LLMs used, the number of in-context examples included, and the evaluation metric.
+With our survey, we hope to give researchers a better estimate of the current progress of goal-oriented prompt engineering in different tasks and also promote a unified evaluation procedure for each task for fair comparison of future methods.
+4 Challenges & Opportunities
+In this section, we discuss the challenges and opportunities of goal-oriented prompt engineering based on our findings.
+Hierarchical Goal Decomposition.
+Existing works only decompose the goal into one layer of sub-goals.
+However, the sub-goal itself can still be complex and multifaceted.
+When a certain sub-goal is complex, LLMs may not achieve it perfectly, and the error will propagated, leading to the failure of the final goal.
+To make sure each sub-goal is simple enough to be handled by LLMs, researchers are now exploring hierarchical decomposition, which further decomposes sub-goals into simpler "sub-sub-goals".
+Li et al.
+(2023) is a pioneering work that constructs a dataset for the script generation task focusing on hierarchical decomposition.
+GITM Zhu et al.
+(2023) employs the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones.
+Though a significant success rate increase by employing hierarchical decomposition is reported, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
+Efficiency.
+The current design of goal-oriented prompt engineering focuses on improving the accuracy of models.
+Nevertheless, efficiency also plays a crucial role in the real-world application of the methods.
+Researchers should consider the trade-off between accuracy and efficiency when necessary.
+We propose to consider two perspectives of efficiency measurement, task-agnostic efficiency and task-specific efficiency.
+Task-agnostic efficiency aims at measuring the efficiency of model targets for any task, for instance, the total number of prompting times before reaching the sub-goal or final goal.
+On the other hand, task-specific efficiency is designed to measure efficiency that appears only in certain tasks.
+As an example, the number of iterations that a certain method prompts LLMs would affect the speed of plan generation in planning tasks.
+Such task-agnostic efficiency is important in scenarios like autonomous driving path planning when we need a quick response for better user experiences.
+On the other hand, the output plan needs to fulfill other criteria, such as minimizing traveling time or distance.
+Here, the time and distance are task-specific efficiency.
+Stage Synergy.
+Overall there is a lack of systematic synergy of stages in any task.
+We found two works that combine their methods, on one stage, with a method, on another stage, in their ablation studies and show performance improvement Chen et al.
+(2022); Wang et al.
+(2023b).
+This suggests the power of synergizing the stages in the goal-oriented framework and careful integration of stages would have the potential for further improvement.
+We take the arithmetic reasoning task as an example.
+To solve a math problem, a naive solution would be to follow Plan-and-solve Wang et al.
+(2023b) to decompose the math problem into a few logical steps.
+Then, following MWP Zhang et al.
+(2023), an operation prediction module can be applied to determine the correct mathematical operations.
+LLMs can select the operands for performing the operation.
+Then, LLMs could call an external calculator to calculate the equation to ensure an accurate result.
+With the result, we could follow Reflexion Shinn et al.
+(2023) to ask LLMs to double-check the intermediate steps again.
+We could further repeat the described process to get multiple sequences of sub-goals and follow MCR Yoran et al.
+(2023) to select valuable sub-goals for more accurate answers.
+Of course, a naive combination of methods from all stages is heavy, requiring additional attention paid to efficiency.
+Additionally, it is also important to consider the between-stages interaction.
+For the same math problem, if we apply a calculator for calculation (action execution) to guarantee a correct answer, we should be careful when designing the result evaluation stage.
+If we use LLMs for result evaluation, the correct result from the calculator may inversely be considered as wrong due to the hallucination problem.
+Broader Applications.
+From our summarization, we observe an absence of study in certain stages for some tasks, such as code generation and recommender system.
+There is potential to improve current methods by studying the unexplored stages.
+In code generation, all of the surveyed methods focus on leveraging sub-goal evaluation to refine the code generation result.
+Additional refinement may be done by incorporating other stages.
+Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to realize the functionality of the sub-function.
+As for the implementation stage of these sub-functions, we could search relevant documents of potential well-written function APIs for reference instead of generating the code fully based on LLMs’ knowledge.
+APIs in external libraries are regularly updated which can perform better than the static knowledge stored in LLMs.
+In recommender systems, works have been proposed for enhancing action execution and sub-goal result evaluation.
+There is an absence of work on action selection.
+Take keeping user retention as the overall goal, one sub-goal might be to make sure that users are recommended with their preferred items.
+There are various actions we could take to predict the preference of a user for a certain item, such as inferring based on the click rate of similar items and inferring based on the watching time of relevant advertisements.
+Careful design on dynamically selecting between available actions based on the characteristics of the user may enable more accurate prediction.
+Moreover, goal-oriented prompt engineering has the potential for broader applications, in for example software development and medical diagnosis, beyond the 11 tasks summarized in this paper.
+For instance, code review is a crucial but cumbersome task in software development.
+It ensures the quality and maintainability of code but requires a reviewer to spend a lot of time doing the systematic examination.
+Automatic code review can help save such significant effort.
+Due to the coding ability shown by recent LLMs, there is potential to surpass existing static analysis tools in automatic code review if equipped with our goal-oriented framework.
+We found one primary study Tufano et al.
+(2024), which focuses on action selection, already shows that reducing the action space by identifying the type of request changes can significantly improve the success rate of ChatGPT.
+Here we give another possibility which covers more stages of our framework.
+With reviewing code as the goal, sub-goals can be documentation quality, function reusability, code style consistency, variable naming readability, and so on.
+Taking variable naming readability as the sub-goal, we can have LLMs select action between checking the self-consistency of the variable names or comparing those names with the pre-defined naming convention (the action selection stage).
+Subsequently, LLMs can execute the action by either calling an external tool for similarity measurement or following given heuristic rules (the action execution stage).
+After going through all the sub-goals (i.e.
+reviewing criteria), LLMs can revise the given code based on reviews from each perspective.
+5 Conclusion
+In this paper, we provide a review of existing research on LLMs prompting from a goal-oriented perspective.
+We explicitly summarize the task-agnostic internal steps involved in designing and refining prompting procedures into five stages, classify the surveyed methods by their targeted tasks, and provide a comprehensive analysis on task coverage and performance.
+Though significant progress has been made in this field, further research is needed to fully unleash the potential by studying hierarchical goal decomposition, the synergy of stages, application to other tasks, and considering efficiency problems.
+We hope this paper can serve as an overview of the current state of goal-oriented prompt engineering and stimulate further research on this important topic.
+Appendix A Peroformance Comparison
+In this section, we compare the performance of surveyed methods under several tasks.
+Due to the different settings in papers, we cannot directly take the reported metric score.
+To address this problem, we first collect scores from those published papers.
+Note that we not only take the score of the proposed methods in those papers but also that of baselines if they appear as one of our surveyed works.
+Once the settings of the same method in different papers have no overlap with each other, we take the score under the setting that has the most similar performance.
+For preprint papers, we take their result once their compared baselines are already recorded and the discrepancy is no more than 3%.
+As a result, the reported performance in our survey is not strictly a fair comparison, and we don’t make conclusions stating which method works best in certain contexts.
+We mainly aim to show that the introduction of goal-oriented prompt design can consistently bring improvement.
+The results are shown in Table , 2, 3, 4, 5, 6, 7, 8.
+Methods VirtualHome
+Zero-shot Planner 46.1
+Re-prompting 41.8
+SALP 49.0
+Table 2: Performance comparison on planning in virtual environment. The evaluation metric is solve rate (%).
+Metric Methods CSQA StrategyQA
+Solve rate CoT 79.0 75.3
+Self-consistency 81.5 81.6
+Plan-and-solve 71.9 65.4
+F1 MCR - 73.6
+Self-consistency - 72.2
+Self-ask - 69.3
+Table 3: Performance comparison on commonsense reasoning under solve rate (%) and F1 (%).
+Methods HumanEval MBPP
+CoT 66.5 48.8
+Self-refine 65.2 48.8
+Reflexion 91.0 77.1
+Self-debug - 75.6
+INTERVENOR 75.6 69.8
+Table 4: Performance comparison on code generation. The evaluation metric is Pass@1 (%).
+Methods Amazon Review Yelp
+CoT 0.7167 0.9794
+ToT 0.7059 0.9766
+Selection-Inference 0.6892 0.9698
+Table 5: Performance comparison on rating prediction in recommender system. The evaluation metric is Mean Absolute Error.
+Methods GSM8K SVAMP ASDiv AQuA AddSub MultiArith
+CoT 56.4 69.9 74.0 38.9 85.3 83.8
+Self-consistency 74.4 86.6 81.9 48.3 93.7 99.3
+PoT 57.0 70.8 - 43.9 85.1 92.2
+Plan-and-solve 59.3 75.7 - 46.0 92.2 91.8
+Table 6: Performance comparison on arithmetic reasoning. The evaluation metric is solve rate (%).
+Methods Symbolic Reasoning Logical Reasoning
+last letter concatenation coinflip PrOntoQA
+CoT 70.4 99.0 87.8
+Self-consistency 73.4 99.5 89.8
+Plan-and-solve 75.2 99.6 -
+RAP - - 94.2
+Table 7: Performance comparison on symbolic reasoning and logical reasoning. The evaluation metric is solve rate (%).
+Metric Methods 2WikiMultihopQA MuSiQue HotpotQA
+Accuracy CoT 29.8 12.6 -
+Least-to-most 29.0 16.8 -
+Self-ask 40.1 16.3 -
+F1 MCR 67.9 - 57.0
+Self-consistency 65.4 - 56.4
+Self-ask 63.8 - 50.2
+DecomP 64.1 29.3 56.4
+Table 8: Performance comparison on multihop question answering under solve rate (%) and F1 (%).
+Application Sub-category Stage Method
+Reasoning Arithmetic Reasoning ①②③④⑤ CoT Wei et al.
 (2022), Self-consistency Wang et al.
@@
 (2023), Plan-and-solve Wang et al.
-(2023a), MWP Zhang et al.
+(2023b), MWP Zhang et al.
 (2023), RAP Hao et al.
-(2023b), SelfCheck Miao et al.
+(2023), SelfCheck Miao et al.
 (2023), Toolformer Schick et al.
@@
 (2022)
-Commonsense Reasoning
-CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023a), Toolformer Schick et al. (2023)
-Symbolic Reasoning
-Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023a)
-Logical Reasoning
-Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023b)
-Planning Virtual Environment ①②④
-Zero-shot Planner Huang et al.
+Commonsense Reasoning CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023b), Toolformer Schick et al. (2023)
+Symbolic Reasoning Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023b)
+Logical Reasoning Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023)
+Planning Virtual Environment ①②④ Zero-shot Planner Huang et al.
 (2022a), Re-prompting Raman et al.
 (2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al.
-(2023b), Reflexion Shinn et al.
+(2023), Reflexion Shinn et al.
 (2023), LLM+P Liu et al.
@@
 (2023), DEPS Wang et al.
-(2023c), GITM Zhu et al.
+(2023d), GITM Zhu et al.
 (2023), Inner Monologue Huang et al.
 (2022b)
-Real Environment
-Inner Monologue Huang et al. (2022b)
-Question Answering Multihop QA ①③④⑤
-DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
-Open domain QA
-Toolformer Schick et al. (2023)
-Code Generation
-Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023)
-Dialogue ①④⑤
-ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
-Recommendation
-Recmind Wang et al. (2023b)
-Table 2: Applicable tasks from existing works.
-The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig.
-1.
-We maintain a list of related datasets and miscellaneous tasks at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering.
-In this section, we provide an overview of the application of goal-oriented prompt engineering on LLMs.
-Specifically, we first introduce the definition of the task and then describe how current methods, following into different stages within our framework, leverage goal-oriented prompt engineering on LLMs to achieve task objectives.
-A summary of the applicable tasks, stages involved, and respective research works are shown in Table 2.
-Note that here we only introduce representative tasks.
-For miscellaneous tasks and related datasets, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering to give more comprehensive information.
-We can see that our framework can be applied to a wide range of tasks.
-Reasoning.
-There are specifically four types of reasoning tasks: arithmetic reasoning, commonsense reasoning, logical reasoning, and symbolic reasoning.
-These reasoning tasks focus on math word problems, world knowledge, logical principles and rules, and manipulation of symbols such as letters, respectively.
-Reasoning involves multiple intermediate steps to obtain a final answer.
-Additionally, to obtain the correct answer the LLM must interpret the question correctly, extract necessary information from the given context, and maintain consistency within its intermediate steps.
-The goal of these tasks can be considered as producing an answer to a given question.
-The intermediate steps can be viewed as sub-goals that are sequentially connected to reach the final goal Wei et al.
-(2022); Wang et al.
-(2023a); Khot et al.
-(2022).
-The sub-goal results from action implementation, which in this case are intermediate conclusions, are often evaluated and revised to avoid error accumulation which may negatively influence the following derivations Madaan et al.
-(2023); Miao et al.
-(2023).
-Alternatively, for certain actions like calculations, we could leverage an external calculator to ensure a precise sub-goal result Schick et al.
-(2023) and employ sketches to help equation formulation (action selection) Zhang et al.
-(2023).
-We can also ask LLMs to explore parallel possible intermediate conclusions to form an overall better solution.
-Planning.
-Given an objective like “Fetch a bottle from the kitchen”, planning involves developing a series of sub-objectives (e.g.
-“Go to the kitchen”, “Find the cupboard”) Liu et al.
-(2023a); Song et al.
-(2023) and reaching sub-objectives by selecting appropriate actions (e.g.
-“Move”, “Search”) Hao et al.
-(2023a); Raman et al.
-(2022) to achieve the overall goal in a virtual or real environment.
-We can see that works focusing on this task do not propose methods for action implementation and valuable sub-goal selection.
-This is because the evaluated environments are relatively straightforward and existing works mainly focus on the accuracy of planning without considering efficiency.
-LLMs often interact with virtual environments through APIs Wang et al.
-(2023c); Zhu et al.
-(2023) or real environments through mechanical devices Huang et al.
-(2022b) to ensure the actions are implemented precisely.
-Question Answering.
-Open domain question answering is a task aimed at answering factual questions without any explicit evidence.
-Multihop QA is an advanced task over open-domain QA.
-The former relies more on world knowledge while the latter requires multi-step reasoning based on world knowledge to answer more complex questions.
-Considering answering complex questions as the goal, LLMs are asked to first decompose them into basic sub-questions Khot et al.
-(2022); Sun et al.
-(2023), gain knowledge for sub-questions Schick et al.
-(2023), and reflect on the correctness of the knowledge Press et al.
-(2022); Shinn et al.
-(2023).
-Sometimes LLMs may ask irrelevant sub-questions or give wrong answers to those sub-questions.
-To avoid their negative influence, there are methods proposed to lead LLMs to select valuable sub-questions from all candidates Yoran et al.
-(2023).
-Code Generation.
-Code generation aims to write a code snippet according to the given description.
-In this task, our goal is to generate code that is both functionally correct and aligns with the description.
-All of the works applied to this task fall into the evaluation of results.
-They leverage the error messages from unit tests Shinn et al.
-(2023); Chen et al.
-(2023) or rely on LLMs themselves Chen et al.
-(2023); Madaan et al.
-(2023) to reflect the validity of generated code snippets and revise them if needed.
-Dialogue.
-Dialogue systems are designed to provide functional service or social support to users via interactions.
-LLMs often play the role of a conversational agent to achieve certain goals (e.g.
-query clarification, offer emotional support).
-To give users positive interaction, LLMs have to correctly understand the context and generate clear and brief responses.
-Thus, some sub-goals are pre-defined according to the goal to ensure effective decomposition Deng et al.
-(2023).
-And LLMs are asked to reflect on their response to generate better content Madaan et al.
-(2023).
-Recommendation.
-The recommender system, which targets providing personalized recommendations to users based on their preferences, is crucial in many application scenarios including social media, e-commerce websites, etc.
-Considering giving ratings for items as the final goal, collecting rating information from various sources are then sub-goals.
-To rely on more valuable sources, LLMs can be asked to compare their credibility and authenticity, and finally take the most valuable source as the overall ratings Wang et al.
-(2023b).
-4 Challenges & Opportunities
-In this section, we discuss the challenges and opportunities of goal-oriented prompt engineering.
-Even if there have been numerous works proposed to facilitate processing at each stage, there remain some potential directions.
-Synergy of stages.
-The synergy of each stage in our framework will further improve the performance of goal-oriented prompt engineering.
-We find that though numerous works are proposed to address different stages in the framework, no one synergizes them all.
-For reasoning tasks, each stage has specially proposed methods, and putting them together is straightforward.
-However, for some tasks like code generation, the vacancy of works for certain stages hinders the integration.
-According to Table 2, there are only works focusing on sub-goal evaluation in code generation.
-To bridge the gap, we have to understand how the code generation process can fit into our framework.
-In fact, programming is a process of breaking down a problem into small, manageable parts and solving them in a logical sequence.
-For example, suppose we are tasked with calculating the average of a set of numbers.
-We can break this problem down into the following steps: (1) Input a set of numbers.
-(2) Calculate the sum of these numbers.
-(3) Determine the count of numbers.
-(4) Use the sum and count to calculate the average.
-Considering these intermediate steps as sub-goals, we decompose the goal of finishing a programming problem into a sequence of sub-goals.
-For other stages like action selection and sub-goal evaluation, we could apply methods that help LLMs efficiently select actions to calculate the summation or reflect whether their intermediate answer is correct.
-Application to other tasks.
-Though existing works have covered up to 11 tasks, goal-oriented prompt engineering has the potential for broader applications.
-For example, apart from implementing functions, code review also plays a crucial role in software development.
-Various review criteria can be taken as sub-goals to ensure the quality of the code, which is our overall goal.
-We may prompt LLMs to design approaches to evaluate different criteria (e.g.
-text-based similarity measurement for coding style evaluation, unit tests generation for functionality verification), which corresponds to the action selection to achieve sub-goals.
-Similar to Program-of-thoughts Chen et al.
-(2022), we can use external interpreters to test the code.
-For some criteria without precise success or failure signals, we can apply sub-goal self-evaluation methods to facilitate better feedback.
-This fitting-in is also suitable for requirement analysis where we also have to think through several perspectives.
-Taking one step further, we think this framework even has the potential to be applied to the whole life cycle of software development.
-Efficiency.
-The works we review in this paper mainly address the accuracy of goal-oriented prompt engineering.
-In other words, they care whether the goal is achieved finally.
-However, efficiency also plays a crucial role in application.
-As we are optimistic about the accuracy when employing methods in each stage of our framework to prompt design, we think efficiency is the next key factor that affects performance.
-For example, in planning tasks, LLMs are expected to form a plan that costs the least energy of human beings or robots.
-And in dialogue systems, users prefer to achieve their goals with as few interaction rounds as possible.
-To evaluate efficiency, we need to either design heuristic rules for automatic measurement or annotate the resource-consuming information by human beings.
-Hierarchical decomposition.
-Existing works decompose the goal into a sub-goal sequence, which can be limiting.
-The method might fall short in adequately addressing the intricacies of each sub-goal, especially in situations where each sub-goal itself is complex and multifaceted.
-Thus, LLMs may not achieve certain sub-goals perfectly, then the error will be accumulated, leading to the failure of overall goal achievement.
-To make sure each sub-goal is simple enough to be handled by LLMs, we introduce the idea of hierarchical decomposition where sub-goals can be decomposed into even more simple sub-goals.
-GITM Zhu et al.
-(2023) could be viewed as employing the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones.
-Though a significant success rate increase by employing hierarchical decomposition is reported, relying on solely a tree may not be able to handle more complex problems, which leaves room for further improvement.
+Real Environment Inner Monologue Huang et al. (2022b)
+Question Answering Multihop QA ①③④⑤ DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
+Open domain QA Toolformer Schick et al. (2023)
+Code Generation Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023), INTERVENOR Wang et al. (2023a)
+Dialogue ①④⑤ ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
+Recommendation ③⑤ Recmind Wang et al. (2023c)
+Table 9: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1.

### v3

@@
 This paper, not only as an overview of current prompt engineering methods, but also aims to highlight the limitation of designing prompts based on an anthropomorphic assumption that expects LLMs to think like humans.
-From our review of 36 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs.
+From our review of 50 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs.
 Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework.
@@
 (2023); Yu et al.
-(2023), reasoning Wei et al.
+(2023a), reasoning Wei et al.
 (2022); Yao et al.
-(2023); Besta et al.
+(2023b); Besta et al.
 (2023), planning Song et al.
@@
 The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into LLMs’ potential Liu et al.
-(2023c).
+(2023d).
 However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al.
@@
 (2023).
-On the other hand, as shown in our literature review of 36 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly.
+On the other hand, as shown in our literature review of 50 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly.
 In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers from LLMs and results in more effective and reliable human-computer interaction.
@@
 Note that stages (2)(3)(4) are taken for all the decomposed sub-goals.
-While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures.
-Some surveys summarize studies that focus on specific tasks or problems.
-Qiao et al.
-(2022); Huang and Chang (2022) focus solely on reasoning tasks and Wang et al.
-(2024) on planning tasks.
-Huang et al.
-(2023) deals with the hallucination problem of LLMs.
-Other studies focus on generally applicable prompting but different sub-fields.
-Amatriain (2024); Sahoo et al.
-(2024) simply listed out recent techniques without systematic categorization.
-Liu et al.
-(2023c) focus on the prompting framework for facilitating the interaction between LLMs with external worlds.
-Liu et al.
-(2023b) focuses on methods for learnable continuous prompt, while our work studies discrete prompt.
-As many of the will-functioned LLMs are not open-sourced (e.g.
-GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
+While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures [^1] We discuss related surveys in Appendix A..
 In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action execution, sub-goal result evaluation, and sub-goal selection, as shown in Fig.
@@
 Finally, we offer four promising future directions, including hierarchical decomposition, efficiency problems, the synergy of stages, and broader applications.
+2 Preliminary
+In this section, we provide a definition of goal-oriented prompt engineering.
+Before define goal-oriented prompting, we can first define standard prompting, the process of prompting LLM $Pr_{LLM}$ for a prediction on a given goal as:
+$Pr(\mathcal{P}|\mathcal{G})=\prod_{i=1}^{|\mathcal{P}|}Pr_{LLM}(p_{i}|\mathcal{G},p_{<i})$ (1)
+By introducing additional variables sub-goal $g$, action $a$, and sub-goal result $r$, we formally define goal-oriented prompt engineering as the application of a multi-stage strategy to guide LLMs in problem-solving by adhering to established human logic, as shown in Fig.
+1.
+In goal-oriented prompt engineering, the overall goal $\mathcal{G}$ is first decomposed into sub-goals $\{g_{1},g_{2},...,g_{n}\}$ (Fig.1(1)), then actions $\{a_{1},a_{2},...,a_{n}\}$ are selected to achieve corresponding sub-goals (Fig.1(2)).
+The final prediction is based on valuable sub-goal results $\{r_{1},r_{2},...,r_{n}\}$ (Fig.1(5)).
+The process could be described as:
+$Pr(\mathcal{P}|\mathcal{G})$ $=\sum_{i=1}^{n}Pr(P|r_{i},a_{i},g_{i},\mathcal{G})$ (2)
+$Pr(r_{i}|a_{i},g_{i},\mathcal{G})Pr(a_{i}|g_{i},\mathcal{G})Pr(g_{i}|\mathcal{G})$
+If LLMs are applied for the implementation of a stage, we could follow Eq.
+1 to get the output.
+We additionally introduce the action execution (Fig.1(3)) and sub-goal evaluation stages (Fig.1(4)) to ensure the correctness of sub-goal results.
 3 Taxonomy of Methods
@@
 With either “Let’s think step by step.” Kojima et al.
-(2022), or intermediate reasoning steps in in-context examples in the prompt, LLMs can imitate the decomposition process and solve the problem step by step.
-Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to a correct output.
+(2022) in the prompt, or intermediate reasoning steps in in-context examples, LLMs can imitate the decomposition process and solve the problem step by step.
+Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence.
 CoT implicitly follows iterative decomposition as LLMs generate tokens in an autoregressive way, with the next sub-goal decided based on previous content.
@@
 DecomP Khot et al.
-(2022) and Successive Prompting Dua et al.
-(2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks.
-The background knowledge required to answer the question is sub-goal to be achieved.
-Different from CoT which may generate sub-questions sequentially in one output, these two works repeatedly prompt LLMs to ask follow-up questions.
+(2022), Successive Prompting Dua et al.
+(2022), and Recursion-of-Thought (RoT) Lee and Kim (2023) are three works that repeatedly prompt LLMs to ask follow-up sub-questions for background knowledge.
 Self-ask Press et al.
-(2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions.
+(2022a) improves efficiency by designing a template for follow-up questions in advance.
 With that template, LLMs can generate all essential questions and their answers by prompting once.
@@
 Plan-and-solve prompting Wang et al.
-(2023b) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output.
+(2023d) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output.
 DEPS Wang et al.
-(2023d) and GITM Zhu et al.
+(2023f) and GITM Zhu et al.
 (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft different items and tools.
@@
 3.2 Action Selection for Attaining Sub-goals
-With sub-goals defined, another important step is action selection, choosing the most effective and valid actions to reach desired outcomes.
+With sub-goals defined, another important step is action selection, choosing effective and valid actions to reach desired outcomes.
 CoT is a naive way for action selection where it is completely decided by LLMs themselves.
-However, due to the hallucination problem, actions generated by LLMs are often invalid.
+However, due to the hallucination problem, actions from LLMs are often invalid.
 In this section, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
@@
 PEARL Sun et al.
-(2023) utilizes constrained action space for question answering over long documents.
-The authors predefine actions like “Finding the definition of A”, “Compare A and B”, and “Summarize A”.
+(2023) utilizes constrained action space like “Finding the definition of A”, “Compare A and B”, and “Summarize A” for question answering over long documents.
 DecomP Khot et al.
 (2022) selects actions from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task.
-SayPlan Rana et al.
-(2023) is for robot planning.
-Given a task instruction, the authors apply semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, then have LLMs plan based on the subgraph.
-In dialogue systems, RLP Fischer (2023) allows the user to set a mental state of LLM, such as beliefs or desires, to instruct action selection.
+In robot planning, given a task instruction, SayPlan Rana et al.
+(2023) applies semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, which makes it easier for LLMs to plan based on the subgraph.
+In dialogue systems, RLP Fischer (2023), SAFARI Wang et al.
+(2023b), and Cue-CoT Wang et al.
+(2023c) set the mental states of the LLM and the user, respectively, such as beliefs or desires, to guide action selection.
 Such a setting improves the context richness, coherence, and interactivity of LLM-based conversational systems.
@@
 Zero-shot planners Huang et al.
-(2022a), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al.
+(2022b), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al.
 (2022) aim to solve agent planning problems in an interactive environment.
@@
 In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result.
-Due to the hallucination problem, the results are sometimes wrong.
+Limited by hallucination, the results are sometimes wrong.
 For example, even if LLMs correctly select multiplication as the action to solve a math problem, the calculation result can be wrong.
-Besides, as Khot et al.
-(2022) claimed, some tasks may not be feasible to solve using only one LLM.
+As Khot et al.
+(2022) claimed, some tasks may not be feasible to solve using only LLMs.
 With such consideration, some works leverage external tools to guarantee a precise sub-goal result.
@@
 (2022); Wang et al.
-(2023c).
+(2023e).
 To answer open-domain questions, DecomP Khot et al.
@@
 Program-of-Thoughts (PoT) Chen et al.
-(2022) first translates reasoning processes into Python codes and then runs them in an interpreter.
-Recmind Wang et al.
-(2023c) uses SQL to search for the history of a user in the database in recommender systems.
+(2022), Faithful CoT Lyu et al.
+(2023), PAL Gao et al.
+(2023), MathPrompter Imani et al.
+(2023), LINC Olausson et al.
+(2023), and Logic-LM Pan et al.
+(2023) translate reasoning processes into executable codes and then run them in an interpreter.
+In recommender systems, Recmind Wang et al.
+(2023e) and InteRecAgent Huang et al.
+(2023b) use SQL to search for the history of a user in the database while DOKE Yao et al.
+(2023a) achieves it through knowledge graphs.
 There are additional studies that aim for multi-tasking.
@@
 Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks.
-Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
+With the development of open-sourced communities, this work can potentially achieve all the sub-goals that can be described in language.
 3.4 Evaluating Sub-goal Results
 Feedback Source Method
-LLM Self-refine Madaan et al. (2023)
-SelfCheck Miao et al. (2023)
-Reflexion Shinn et al. (2023)
-Environment SayPlan Rana et al. (2023)
-Re-prompt Raman et al. (2022)
-Inner Monologue Huang et al. (2022b)
-GITM  Zhu et al. (2023)
-DEPS  Wang et al. (2023d)
-VLM DEPS  Wang et al. (2023d)
-LLM-Planner Song et al. (2023)
-Inner Monologue Huang et al. (2022b)
-Program Executor Self-debug Chen et al. (2023)
-INTERVENOR Wang et al. (2023a)
+LLM Self-refine Madaan et al. (2023), SelfCheck Miao et al. (2023), Reflexion Shinn et al. (2023), REFINER Paul et al. (2024)
+Environment SayPlan Rana et al. (2023), Re-prompt Raman et al. (2022), Inner Monologue Huang et al. (2022c), GITM Zhu et al. (2023), DEPS Wang et al. (2023f)
+VLM DEPS  Wang et al. (2023f), LLM-Planner Song et al. (2023), Inner Monologue Huang et al. (2022c)
+Interpreter Self-debug Chen et al. (2023), INTERVENOR Wang et al. (2023a), CRITIC Gou et al. (2023), MAF Nathani et al. (2023)
 Heuristic Rule Reflexion Shinn et al. (2023)
-Human Inner Monologue Huang et al. (2022b)
-DEPS  Wang et al. (2023d)
+Human Inner Monologue Huang et al. (2022c), DEPS Wang et al. (2023f)
+Search Engine CRITIC Gou et al. (2023), Verify-and-Edit Zhao et al. (2023)
 Table 1: Feedback source used by existing works.
-In CoT, there is no evaluation process for the sub-goal results.
-The results generated by LLMs are assumed to be correct by default.
-Answering by LLMs, sub-goal results also suffer from the hallucination problem.
-In addition, errors may propagate along the pipeline and result in huge deviations from the correct answer.
-Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time.
+In CoT, the results generated by LLMs are assumed to be correct hence there is no evaluation process for the sub-goal results.
+The negative impact of hallucination becomes more severe for sub-goal results because errors may propagate along the pipeline and result in huge deviations from the correct answer.
+Thus, it is necessary to get feedback for sub-goal results at each step and correct them in time through resampling.
+The types of feedback sources are summarized in Table 1.
 Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well.
-The types of feedback source and their related works are summarized in Table 1.
 Figure 2: Illustration of approaches for valuable sub-goal selection.
@@
 (g) GoT models all sub-goals in a graph structure space.
-Evaluate with LLMs.
+Feedback from LLMs.
 Self-refine Madaan et al.
-(2023) is the first work addressing the evaluation of sub-goal results.
+(2023); Huang et al.
+(2022a) is the first work addressing the evaluation of sub-goal results.
 After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output.
@@
 The refinement process iterates with all of the past feedback and refinement appended to the prompt.
-For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified.
 SelfCheck Miao et al.
@@
 In Reflexion Shinn et al.
-(2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result.
-Evaluate with external evaluators.
-In contrast to self-evaluation, we could seek the help of external evaluators for providing feedback, which is similar to but different from the idea of applying external tools for action execution.
+(2023) and REFINER Paul et al.
+(2024), the authors try using another instantiation of an LLM to obtain feedback for the sub-goal result.
+Feedback from external tools.
+In contrast to self-evaluation, we could seek the help of external evaluators to provide feedback.
 Some methods Chen et al.
@@
 Self-debug Chen et al.
+(2023), MAF Nathani et al.
 (2023) and INTERVENOR Wang et al.
@@
 The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection.
-This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback.
+This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback Gou et al.
+(2023).
 SayPlan Rana et al.
@@
 GITM Zhu et al.
-(2023), a method proposed for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation.
-All of the information can be obtained from Minecraft API returns.
-The quality of error messages (e.g.
-detail) is influential to the evaluating process, but sometimes it is hard to define error messages in advance.
+(2023) designs a structured feedback template to get detailed knowledge of the current situation and all of the information can be obtained from environment API returns.
+However, sometimes it is hard to define error messages in advance.
 To further address it, LLM-Planner Song et al.
-(2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed.
+(2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment.
 Inner Monologue Huang et al.
-(2022b) is a combination of SayPlan and LLM-Planner, where VLMs additionally perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information.
-Similar to Inner Monologue, DEPS Wang et al.
-(2023d), designed for Minecraft, also gets feedback from both VLMs and environments.
-DEPS further introduces a step to ask LLMs to explain the reason according to the feedback.
+(2022c) and DEPS Wang et al.
+(2023f) are combinations of SayPlan and LLM-Planner, where VLMs additionally perform visual question answering to enable a finer-grained human-computer interaction for valuable information.
 Reflexion Shinn et al.
-(2023) proposes various evaluators to handle different tasks.
-It has reward functions for reasoning tasks based on exact match grading and heuristic functions tailored to specific criteria for decision-making tasks.
-We refer to them as heuristic rules.
-Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine.
-Instead of executing the action to get the sub-goal result again, these methods restart from action selection.
+(2023) proposes various heuristic functions to score different tasks (e.g.
+exact match for reasoning).
+CRITIC Gou et al.
+(2023) and Verify-and-Edit Zhao et al.
+(2023) employ search engines to obtain true information as feedback from the web.
 3.5 Further Selection of Valuable Sub-goals
-In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process.
-Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant.
-To address this problem, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions.
-Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement.
-In this section, we introduce methods for such sub-goal postprocessing, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al.
-(2023) and its variants, and finally Graph of thoughts (GoT) Besta et al.
+The generated sub-goals in stage (1) may be wrong or irrelevant to overall goal achievement.
+To address this problem, we could ask LLMs to explore several sub-goals in each step or even several sub-goal sequences and then only select those valuable candidates.
+In this section, we introduce methods for such sub-goal post-processing, which can be divided into three stages of development, CoT and its variants, Tree of thoughts (ToT) Yao et al.
+(2023b) and its variants, and finally Graph of thoughts (GoT) Besta et al.
 (2023).
@@
 We consider CoT as simply selecting all sub-goals since LLMs have access to all of the sub-goals during the generation process, which makes it prone to irrelevant sub-goals or wrong sub-goal results.
-Four works are proposed to improve the robustness of CoT through sub-goal postprocess.
+Four works are proposed to improve the robustness of CoT.
 Self-consistency Wang et al.
-(2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences.
+(2022) prompts LLMs several times via CoT and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences.
 The result from the chosen sequence is considered the final result.
@@
 Recmind Wang et al.
-(2023c) and Selection-Inference Creswell et al.
+(2023e) and Selection-Inference Creswell et al.
 (2022) allow multiple sub-goal candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
@@
 ToT Yao et al.
-(2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure.
+(2023b) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure.
 To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search.
@@
 (2023) and GDP-Zero Yu et al.
-(2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS).
+(2023a) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS).
 Compared with ToT, MCTS has an additional back-propagation step, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence.
@@
 For more comprehensive information, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering..
-The 36 representative works in our survey cover 11 tasks, including arithmetic reasoning (11 works), commonsense reasoning (5 works), symbolic reasoning (2 works), logical reasoning (3 works), planning in virtual environment (10 works), planning in real environment (1 work), multihop question answering (5 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (3 works), and recommender systems (1 work).
-We could see that 60% of the works focus on reasoning, 31% of the works are for planning and the remaining 9% of the works are for other tasks.
-For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 23% equally.
+The 50 representative works in our survey cover 11 tasks, including arithmetic reasoning (18 works), commonsense reasoning (5 works), symbolic reasoning (3 works), logical reasoning (5 works), planning in virtual/real environment (12 works), multihop question answering (8 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (6 works), and recommender systems (3 works).
+We could see that 46% of the works focus on reasoning, 24% of the works are for planning and the remaining 30% of the works are for other tasks.
+For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 27% equally.
 There is potential for exploring the synergy of stages, which we will discuss in Section 5.
-For planning tasks, goal decomposition shares 31%, action selection shares 23%, and sub-goal result evaluation shares the rest 46%.
+For planning tasks, goal decomposition and action selection share 27%, and sub-goal result evaluation shares the rest 46%.
 No surveyed work for planning tasks involves the action execution stage and valuable sub-goal selection stage.
@@
 Actions are either easy or can be assured to be perfectly executed.
-Nonetheless, as the field develops, we can expect more complicated environments when actions cost differently or cannot be perfectly executed.
-Under such scenarios, there should be attention paid to action execution and valuable sub-goal selection stage.
+Nonetheless, as the field develops, we can expect more complicated environments where there should be attention paid to action execution and valuable sub-goal selection stage.
 For those minority tasks like code generation and recommendation, most of the stages remain unexplored.
@@
 Performance.
-Furthermore, We compare the performance of the surveyed methods in Appendix A (Table 2, 3, 4, 5, 6, 7, 8).
-Compared with vanilla CoT, on average goal decomposition (Plan-and-solve) improves by 6.1%, action selection (PoT) improves by 2.9%, and valuable sub-goal selection (Self-consistency) improves by 22.7% in arithmetic reasoning, as shown in Table 6.
+Furthermore, We compare the performance of the surveyed methods in Appendix B (Table 2, 3, 4, 5, 6, 7, 8, 9), and there are some noteworthy findings.
+In arithmetic reasoning, compared with standard I-O prompting, on average goal decomposition (CoT) improves by 22.6%, sub-goal evaluation (Self-refine Madaan et al.
+(2023)) improves by 21.1%, and valuable sub-goal selection (Self-consistency Wang et al.
+(2022)) improves by 32.5%.
+While valuable sub-goal selection improves the most in arithmetic reasoning, the improvement brought by goal decomposition outperforms sub-goal selection in symbolic reasoning, which indicates that the variations of improvement have relations to task characteristics.
 Besides, we observe a gradual increase in performance when integrating additional stages into the method.
-When combined with Self-consistency, Plan-and-solve can be improved by 11.3% and PoT can be improved by 5.2% on average across several arithmetic reasoning datasets, which indicates that stage synergy may bring further improvement.
-During data collection, we find it hard to compare between methods due to the variety of LLMs used, the number of in-context examples included, and the evaluation metric.
-With our survey, we hope to give researchers a better estimate of the current progress of goal-oriented prompt engineering in different tasks and also promote a unified evaluation procedure for each task for fair comparison of future methods.
+PoT Chen et al.
+(2022) can be considered as a combination of CoT and action execution, it further improves CoT by 14.7% in arithmetic reasoning.
+Similarly, Sayplan Rana et al.
+(2023) in planning tasks, equipped with goal decomposition, action selection, and sub-goal evaluation, outperforms solely sub-goal evaluation (LLM-Planner Song et al.
+(2023)) by 73.3% in excitability.
+Thus, we could expect further improvement from stage synergy.
+During data collection, we find it hard to compare between methods due to the variety of prompt templates, the number of in-context examples included, and the decoding strategies, even if the methods are applied to the same LLM in the same zero-/few-shot setting.
+With our survey, we hope to promote a unified evaluation procedure for each task for fair comparison of future methods.
 5 Challenges & Opportunities
@@
 Hierarchical Goal Decomposition.
-Existing works only decompose the goal into one layer of sub-goals.
-However, the sub-goal itself can still be complex and multifaceted.
-When a certain sub-goal is complex, LLMs may not achieve it perfectly, and the error will propagated, leading to the failure of the final goal.
+Existing works only decompose the goal into one layer of sub-goals, ignoring the fact that the sub-goal itself can still be complex and multifaceted.
 To make sure each sub-goal is simple enough to be handled by LLMs, researchers are now exploring hierarchical decomposition, which further decomposes sub-goals into simpler "sub-sub-goals".
@@
 GITM Zhu et al.
-(2023) employs the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones.
-Though a significant success rate increase by employing hierarchical decomposition is reported, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
+(2023) employs the idea of hierarchical decomposition by utilizing a pre-defined sub-goal tree to connect a sub-goal with even simpler ones.
+Though a significant success rate increase by employing hierarchical decomposition is reported, a pre-defined tree is not always accessible.
+Besides, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
 Efficiency.
@@
 Overall there is a lack of systematic synergy of stages in any task.
-We found two works that combine their methods, on one stage, with a method, on another stage, in their ablation studies and show performance improvement Chen et al.
-(2022); Wang et al.
-(2023b).
+We found two works that combine their methods, on one stage, with a method, on another stage, and show performance improvement Rana et al.
+(2023); Chen et al.
+(2022).
 This suggests the power of synergizing the stages in the goal-oriented framework and careful integration of stages would have the potential for further improvement.
@@
 To solve a math problem, a naive solution would be to follow Plan-and-solve Wang et al.
-(2023b) to decompose the math problem into a few logical steps.
+(2023d) to decompose the math problem into a few logical steps.
 Then, following MWP Zhang et al.
@@
 We could further repeat the described process to get multiple sequences of sub-goals and follow MCR Yoran et al.
-(2023) to select valuable sub-goals for more accurate answers.
+(2023) to remove less valuable equations for more accurate answers.
 Of course, a naive combination of methods from all stages is heavy, requiring additional attention paid to efficiency.
@@
 There is potential to improve current methods by studying the unexplored stages.
+Moreover, goal-oriented prompt engineering has the potential for broader applications beyond the 11 tasks summarized in this paper.
 In code generation, all of the surveyed methods focus on leveraging sub-goal evaluation to refine the code generation result.
 Additional refinement may be done by incorporating other stages.
-Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to realize the functionality of the sub-function.
+Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to implement them.
 As for the implementation stage of these sub-functions, we could search relevant documents of potential well-written function APIs for reference instead of generating the code fully based on LLMs’ knowledge.
@@
 Careful design on dynamically selecting between available actions based on the characteristics of the user may enable more accurate prediction.
-Moreover, goal-oriented prompt engineering has the potential for broader applications, in for example software development and medical diagnosis, beyond the 11 tasks summarized in this paper.
-For instance, code review is a crucial but cumbersome task in software development.
+For tasks beyond the scope of this paper, here we take code review as an example.
+Code review is a crucial but cumbersome task in software development.
 It ensures the quality and maintainability of code but requires a reviewer to spend a lot of time doing the systematic examination.
@@
 We hope this paper can serve as an overview of the current state of goal-oriented prompt engineering and stimulate further research on this important topic.
-Appendix A Peroformance Comparison
-In this section, we compare the performance of surveyed methods under several tasks.
-Due to the different settings in papers, we cannot directly take the reported metric score.
-To address this problem, we first collect scores from those published papers.
-Note that we not only take the score of the proposed methods in those papers but also that of baselines if they appear as one of our surveyed works.
-Once the settings of the same method in different papers have no overlap with each other, we take the score under the setting that has the most similar performance.
-For preprint papers, we take their result once their compared baselines are already recorded and the discrepancy is no more than 3%.
-As a result, the reported performance in our survey is not strictly a fair comparison, and we don’t make conclusions stating which method works best in certain contexts.
+Appendix A Related Surveys
+While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures.
+Some surveys summarize studies that focus on specific tasks or problems.
+Qiao et al.
+(2022); Huang and Chang (2022); Lu et al.
+(2023); Yang et al.
+(2023) focus solely on reasoning tasks and Wang et al.
+(2024) on planning tasks.
+Huang et al.
+(2023a) deals with the hallucination problem of LLMs.
+Other studies focus on generally applicable prompting but different sub-fields.
+Amatriain (2024); Sahoo et al.
+(2024) simply listed out recent techniques without systematic categorization.
+Yu et al.
+(2023b); Chu et al.
+(2023) study CoT and its XoT variants.
+Liu et al.
+(2023d) focus on the prompting framework for facilitating the interaction between LLMs with external worlds.
+Liu et al.
+(2023c) focuses on methods for learnable continuous prompt, while our work studies discrete prompt.
+As many of the will-functioned LLMs are not open-sourced (e.g.
+GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
+Appendix B Performance Comparison
+In this section, we compare the performance of surveyed methods under several tasks, including arithmetic reasoning (Table 2), commonsense reasoning (Table 3), symbolic reasoning (Table 4), logical reasoning (Table 5), planning (Table 6), multi-hop question answering (Table 7), code generation (Table 8), and recommender systems (Table 9).
+The empirical results are collected from corresponding papers or other papers where they serve as baselines.
+Here, the benchmarks we chose are the ones on which the majority of methods are evaluated.
+For a full list of benchmarks, we maintain it at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering.
+Note that due to variations in experimental settings like prompt template, decoding strategy, and numbers of in-context examples, even the methods that share the same zero-/few-shot setting and LLM may not be fairly comparable.
+Therefore, these tables only provide a rough trend of performance, and we don’t make conclusions stating which method works best in certain contexts.
 We mainly aim to show that the introduction of goal-oriented prompt design can consistently bring improvement.
-The results are shown in Table , 2, 3, 4, 5, 6, 7, 8.
-Methods VirtualHome
-Zero-shot Planner 46.1
-Re-prompting 41.8
-SALP 49.0
-Table 2: Performance comparison on planning in virtual environment. The evaluation metric is solve rate (%).
-Metric Methods CSQA StrategyQA
-Solve rate CoT 79.0 75.3
-Self-consistency 81.5 81.6
-Plan-and-solve 71.9 65.4
-F1 MCR - 73.6
-Self-consistency - 72.2
-Self-ask - 69.3
+Methods Stage Setting Model GSM8K SVAMP AQuA AddSub MultiArith
+Cobbe et al. (2021) Patel et al. (2021) Ling et al. (2017) Hosseini et al. (2014) Roy and Roth (2016)
+Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 56.4 69.9 38.9 85.3 83.8
+Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 58.4 80.3 48.4 91.6 93.6
+Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 59.3 75.7 46.0 92.2 91.8
+Least-to-most Zhou et al. (2022) few-shot code-davinci-002 38.3 80.3 40.6 - 74.0
+PoT Chen et al. (2022) zero-shot text-davinci-003 57.0 70.8 43.9 85.1 92.2
+Toolformer Schick et al. (2023) zero-shot GPT-J - 29.4 - - -
+PoT Chen et al. (2022) few-shot code-davinci-002 71.6 85.2 54.1 - -
+PAL Gao et al. (2023) few-shot code-davinci-002 72.0 79.4 - 92.5 99.2
+Faithful CoT Lyu et al. (2023) few-shot code-davinci-002 72.3 83.4 47.2 - 98.8
+RAP Hao et al. (2023) few-shot LLaMA-33B 40.0 - - - -
+Self-refine Madaan et al. (2023) few-shot text-davinci-003 64.1 67.6 - - -
+Self-consistency Wang et al. (2022) few-shot code-davinci-002 78.0 86.8 52.0 91.6 100.0
+Table 2: Performance comparison on arithmetic reasoning. The evaluation metric is solve rate (%).
+Metric Methods Stage Setting Model CSQA StrategyQA
+Talmor et al. (2018) Geva et al. (2021)
+Solve rate Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 65.2 63.8
+Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 71.9 65.4
+Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 78.3 71.2
+Few-shot CoT Wei et al. (2022) few-shot code-davinci-002 79.0 73.4
+Self-consistency Wang et al. (2022) few-shot code-davinci-002 81.5 79.8
+F1 CoT Wei et al. (2022) few-shot code-davinci-002 - 70.0
+Self-ask Press et al. (2022a) few-shot code-davinci-002 - 69.3
+MCR Yoran et al. (2023) few-shot code-davinci-002 - 73.6
+Self-consistency Wang et al. (2022) few-shot code-davinci-002 - 72.2
 Table 3: Performance comparison on commonsense reasoning under solve rate (%) and F1 (%).
-Methods HumanEval MBPP
-CoT 66.5 48.8
-Self-refine 65.2 48.8
-Reflexion 91.0 77.1
-Self-debug - 75.6
-INTERVENOR 75.6 69.8
+F1 is computed by treating prediction and ground truth answers as bags of tokens and computing their precision and recall.
+Methods Stage Setting Model last letter concatenation coinflip
+Wei et al. (2022) Wei et al. (2022)
+I-O Prompting - few-shot text-davinci-002 5.8 49.0
+Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 64.8 96.8
+Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 75.2 99.6
+Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 70.6 100.0
+Few-shot CoT Wei et al. (2022) few-shot code-davinci-002 70.4 99.0
+Least-to-most Zhou et al. (2022) few-shot code-davinci-002 94.0 -
+Self-consistency Wang et al. (2022) few-shot code-davinci-002 73.4 99.5
+Table 4: Performance comparison on symbolic reasoning. The evaluation metric is solve rate (%).
+Methods Stage Setting Model FOLIO ProofWriter
+Han et al. (2022) Tafjord et al. (2020)
+I-O Prompting - zero-shot gpt-3.5-turbo 48.4 36.4
+CoT Wei et al. (2022) few-shot gpt-3.5-turbo 54.9 43.6
+Logic-LM Pan et al. (2023) few-shot gpt-3.5-turbo 62.7 58.3
+LINC Olausson et al. (2023) few-shot gpt-3.5-turbo 62.6 96.4
+LINC Olausson et al. (2023) few-shot GPT-4 72.5 98.3
+Table 5: Performance comparison on logical reasoning. The evaluation metric is solve rate (%).
+Dataset Methods Stage Setting Model Evaluation Metric
+Executability Correctness
+Home+Office LLM+P Liu et al. (2023a) few-shot GPT-4 0.0 33.3
+LLM-Planner Song et al. (2023) few-shot GPT-4 13.3 66.7
+Rana et al. (2023) SayPlan Rana et al. (2023) ①②④ few-shot GPT-4 86.6 73.3
+VirtualHome Zero-shot Planner Huang et al. (2022b) few-shot GPT2-large 16.4 33.3
+Zero-shot Planner Huang et al. (2022b) few-shot Codex 78.4 46.1
+Puig et al. (2018) SALP Gramopadhye and Szafir (2023) few-shot GPT2-large 50.8 49.0
+Re-prompting Raman et al. (2022) few-shot Codex 98.9 41.8
+Table 6: Performance comparison on planning in virtual environment.
+The evaluation metrics are executability (%) and correctness (%).
+Executability measures if the generated actions can be correctly parsed in the environment, while Correctness is a human-annotated metric that assesses whether the given steps could accomplish the given task.
+Metric Methods Stage Setting Model 2WikiMultihopQA Bamboogle HotpotQA
+Ho et al. (2020) Press et al. (2022b) Yang et al. (2018)
+Acc. I-O Prompting - few-shot Codex 25.4 17.6 -
+CoT Wei et al. (2022) few-shot Codex 29.8 46.4 -
+Least-to-most Zhou et al. (2022) few-shot Codex 29.0 - -
+Self-ask Press et al. (2022a) few-shot Codex 40.1 60.0 -
+F1 CoT Wei et al. (2022) few-shot Codex 67.2 64.7 56.4
+Self-ask Press et al. (2022a) few-shot Codex 63.8 64.6 50.2
+DecomP Khot et al. (2022) zero-shot Codex 64.1 25.4 49.9
+MCR Yoran et al. (2023) few-shot Codex 67.9 66.5 57.0
+Self-consistency Wang et al. (2022) few-shot Codex 65.4 65.0 56.4
+Table 7: Performance comparison on multihop question answering under accuracy (%) and F1 (%).
+F1 is computed by treating prediction and ground truth answers as bags of tokens and computing their precision and recall.
+Methods Stage Setting Model HumanEval MBPP
+Chen et al. (2021) Austin et al. (2021)
+Zero-shot I-O Prompting - zero-shot gpt-3.5-turbo 62.2 41.6
+Few-shot I-O Prompting - few-shot gpt-3.5-turbo 65.2 40.6
+CoT Wei et al. (2022) zero-shot gpt-3.5-turbo 66.5 48.8
+Self-refine Madaan et al. (2023) zero-shot gpt-3.5-turbo 65.2 48.8
+INTERVENOR Wang et al. (2023a) zero-shot gpt-3.5-turbo 75.6 69.8
+Reflexion Shinn et al. (2023) zero-shot GPT-4 91.0 77.1
+Self-debug Chen et al. (2023) few-shot code-davinci-002 - 75.6
 Table 8: Performance comparison on code generation. The evaluation metric is Pass@1 (%).
-Methods Amazon Review Yelp
-CoT 0.7167 0.9794
-ToT 0.7059 0.9766
-Selection-Inference 0.6892 0.9698
-Table 5: Performance comparison on rating prediction in recommender system. The evaluation metric is Mean Absolute Error.
-Methods GSM8K SVAMP ASDiv AQuA AddSub MultiArith
-CoT 56.4 69.9 74.0 38.9 85.3 83.8
-Self-consistency 74.4 86.6 81.9 48.3 93.7 99.3
-PoT 57.0 70.8 - 43.9 85.1 92.2
-Plan-and-solve 59.3 75.7 - 46.0 92.2 91.8
-Table 6: Performance comparison on arithmetic reasoning. The evaluation metric is solve rate (%).
-Methods Symbolic Reasoning Logical Reasoning
-last letter concatenation coinflip PrOntoQA
-CoT 70.4 99.0 87.8
-Self-consistency 73.4 99.5 89.8
-Plan-and-solve 75.2 99.6 -
-RAP - - 94.2
-Table 7: Performance comparison on symbolic reasoning and logical reasoning. The evaluation metric is solve rate (%).
-Metric Methods 2WikiMultihopQA MuSiQue HotpotQA
-Accuracy CoT 29.8 12.6 -
-Least-to-most 29.0 16.8 -
-Self-ask 40.1 16.3 -
-F1 MCR 67.9 - 57.0
-Self-consistency 65.4 - 56.4
-Self-ask 63.8 - 50.2
-DecomP 64.1 29.3 56.4
-Table 8: Performance comparison on multihop question answering under solve rate (%) and F1 (%).
+Methods Stage Setting Model Amazon Review Yelp
+Ni et al. (2019) Geng et al. (2022)
+I-O Prompting Liu et al. (2023b) - zero-shot gpt-3.5-turbo 1.1897 1.2359
+CoT Wei et al. (2022) zero-shot gpt-3.5-turbo 0.8612 1.1673
+Selection-Inference Creswell et al. (2022) zero-shot gpt-3.5-turbo 0.7883 1.0009
+I-O Prompting Liu et al. (2023b) - few-shot gpt-3.5-turbo 0.7327 1.0016
+CoT Wei et al. (2022) few-shot gpt-3.5-turbo 0.7167 0.9794
+ToT Yao et al. (2023b) few-shot gpt-3.5-turbo 0.7059 0.9766
+Selection-Inference Creswell et al. (2022) few-shot gpt-3.5-turbo 0.6892 0.9698
+Table 9: Performance comparison on rating prediction in recommender system.
+The evaluation metric is Mean Absolute Error.
+Here I-O prompting means that LLMs are directly prompted for the final prediction.
 Application Sub-category Stage Method
-Reasoning Arithmetic Reasoning ①②③④⑤ CoT Wei et al.
-(2022), Self-consistency Wang et al.
-(2022), Least-to-most Zhou et al.
-(2022), PoT Chen et al.
-(2022), Self-refine Madaan et al.
-(2023), Plan-and-solve Wang et al.
-(2023b), MWP Zhang et al.
+Reasoning Arithmetic CoT Wei et al. (2022), Least-to-most Zhou et al. (2022), Plan-and-solve Wang et al. (2023d), Successive Prompt Dua et al. (2022), RoT Lee and Kim (2023)
+MWP Zhang et al. (2023)
+PoT Chen et al. (2022), Toolformer Schick et al. (2023), Faithful CoT Lyu et al. (2023), PAL Gao et al. (2023), MathPrompter Imani et al. (2023)
+Self-refine Madaan et al. (2023), RAP Hao et al. (2023), SelfCheck Miao et al. (2023), REFINER Paul et al. (2024), CRITIC Gou et al. (2023), MAF Nathani et al. (2023)
+Self-consistency Wang et al. (2022)
+Commonsense CoT Wei et al. (2022), Plan-and-solve Wang et al. (2023d)
+Toolformer Schick et al. (2023)
+Self-consistency Wang et al. (2022), MCR Yoran et al. (2023)
+Symbolic Plan-and-solve Wang et al. (2023d)
+PAL Gao et al. (2023)
+Self-consistency Wang et al. (2022)
+Logical DecomP Khot et al. (2022)
+LINC Olausson et al. (2023), Logic-LM Pan et al. (2023)
+RAP Hao et al. (2023)
+Selection-Inference Creswell et al. (2022)
+Planning Virtual Env LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023f), GITM Zhu et al. (2023)
+Zero-shot Planner Huang et al. (2022b), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), SayPlan Rana et al. (2023),
+Re-prompting Raman et al.
 (2022), RAP Hao et al.
-(2023), SelfCheck Miao et al.
-(2023), Toolformer Schick et al.
-(2023), Successive Prompt Dua et al.
+(2023), Reflexion Shinn et al.
+(2023), SayPlan Rana et al.
+(2023), DEPS Wang et al.
+(2023f), GITM Zhu et al.
+(2023), Inner Monologue Huang et al.
+(2022c), LLM-Planner Song et al.
 (2023)
-Commonsense Reasoning CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023b), Toolformer Schick et al. (2023)
-Symbolic Reasoning Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023b)
-Logical Reasoning Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023)
-Planning Virtual Environment ①②④ Zero-shot Planner Huang et al.
-(2022a), Re-prompting Raman et al.
-(2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al.
-(2023), Reflexion Shinn et al.
-(2023), LLM+P Liu et al.
-(2023a), SayPlan Rana et al.
-(2023), DEPS Wang et al.
-(2023d), GITM Zhu et al.
-(2023), Inner Monologue Huang et al.
-(2022b)
-Real Environment Inner Monologue Huang et al. (2022b)
-Question Answering Multihop QA ①③④⑤ DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
+Real Env Inner Monologue Huang et al. (2022c)
+Question Answering Multihop QA DecomP Khot et al. (2022), Self-ask Press et al. (2022a)
+PEARL Sun et al. (2023)
+Faithful CoT Lyu et al. (2023)
+Reflexion Shinn et al. (2023), CRITIC Gou et al. (2023), Verify-and-Edit Zhao et al. (2023)
+MCR Yoran et al. (2023)
 Open domain QA Toolformer Schick et al. (2023)
 Code Generation Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023), INTERVENOR Wang et al. (2023a)
-Dialogue ①④⑤ ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
-Recommendation ③⑤ Recmind Wang et al. (2023c)
-Table 9: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1.
+Dialogue ProCoT Deng et al. (2023), RLP Fischer (2023), Cue-CoT Wang et al. (2023c), SAFARI Wang et al. (2023b)
+Self-refine Madaan et al. (2023)
+GDP-Zero Yu et al. (2023a),
+Recommendation Recmind Wang et al. (2023e), DOKE Yao et al. (2023a), InteRecAgent Huang et al. (2023b)
+Recmind Wang et al. (2023e)
+Table 10: Applicable tasks from existing works.
+The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig.
+1.
+Note that the methods in the table are not exclusively applicable to the tasks listed in the table to which they belong.
+We list them based on the evaluated tasks in their original papers.